### v2

# Enhancing Generative Auto-bidding with Offline Reward Evaluation and Policy Search

Abstract
Auto-bidding serves as a critical tool for advertisers to improve their advertising performance. Recent progress has demonstrated that AI-Generated Bidding (AIGB), which learns a conditional generative planner from offline data, achieves superior performance compared to typical offline reinforcement learning (RL)-based auto-bidding methods. However, existing AIGB methods still face a performance bottleneck due to their inherent inability to explore beyond the static offline dataset. To address this, we propose AIGB-Pearl (Planning with EvaluAtor via RL), a novel method that integrates generative planning and policy optimization. The core of AIGB-Pearl lies in constructing a trajectory evaluator for scoring generation quality and designing a provably sound KL-Lipschitz-constrained score maximization scheme to ensure safe and efficient generalization beyond the offline dataset. A practical algorithm incorporating the synchronous coupling technique is further devised to ensure the model regularity required by the proposed scheme. Extensive experiments on both simulated and real-world advertising systems demonstrate the state-of-the-art performance of our approach.
1 Introduction
The increasing demand for commercial digitalization has facilitated the development of the auto-bidding technique in online advertising. Distinguished from traditional manual bidding products, auto-bidding provides advertisers with an efficient and flexible scheme to automatically optimize bids in dynamic and competitive environments (Balseiro et al., 2021a; Deng et al., 2021; Balseiro et al., 2021b). Technically, auto-bidding can be viewed as an offline sequential decision-making problem, which aims to maximize the advertising performance over the bidding episode, relying solely on a static offline dataset due to system limitations (Mou et al., 2022), such as safety concerns.
As a standard approach to offline decision-making problems, offline reinforcement learning (RL) (Kumar et al., 2020) is widely adopted to solve the auto-bidding problem. By employing conservative policy search schemes, offline RL mitigates the infamous out-of-distribution (OOD) problem (Fujimoto et al., 2019), enabling reliable generalization beyond the offline dataset. However, due to the employment of bootstrapped value estimations, offline RL methods typically suffer from the training instability issue (Peng et al., 2024), risking policy performance degradation and instability.
Recent advances in generative models shed new light on offline decision-making problems (Zhu et al., 2023; Kang et al., 2023). Specifically, AI-generated bidding (AIGB) models the auto-bidding as a trajectory generation task and employs a generative model to approximate the conditional trajectory distribution of the offline dataset (Guo et al., 2024). AIGB avoids the requirement for bootstrapping and exhibits more stable training and superior performance. However, the modeling approach in AIGB does not explicitly align with the performance optimization objective of the auto-bidding problem, preventing it from actively improving toward higher performance. As a result, AIGB primarily imitates well-behaved trajectories from the offline dataset, lacking the ability to explore higher-quality trajectories beyond the offline data (Ajay et al., 2023).
Hence, there arises a question: built on AIGB, the latest state-of-the-art auto-bidding method, can we devise a plausible scheme to involve policy optimization in its generative model? To this end, a natural idea is to integrate offline RL methods into AIGB. However, it is nontrivial to implement in the auto-bidding problem since (i) there is a lack of reward signals in AIGB to guide the generative model. Specifically, the generation quality of the generative model remains unknown during training, making it infeasible to explore new trajectories beyond the offline dataset; (ii) no dedicated offline RL algorithm exists for AIGB. In particular, theoretical analysis that guarantees safe generalization and mitigates OOD issues for generative models in auto-bidding remains largely unexplored.
To address these critical challenges, we propose AIGB-Pearl (Planning with EvaluAtor via RL), an RL-enhanced version of AIGB that learns a trajectory evaluator to score generation quality and drive exploration through continuous interaction with the generative model in AIGB via RL. The evaluator is trained through supervised learning, leveraging the offline dataset. Notably, to mitigate the OOD problem and ensure reliable usage of the evaluator, we first examine the theoretical upper bound on the evaluator’s bias. Then, guided by this analysis, we establish a KL-Lipschitz-constrained score maximization objective with a provable sub-optimality bound, enabling safe and effective exploration for the generative model beyond the offline data. Moreover, to perform constrained score maximization, we design a practical algorithm incorporating the synchronous coupling technique, which helps ensure the Lipschitz requirement of the generative model. Besides, we note that AIGB-Pearl operates without the need for bootstrapping, leading to greater training stability compared to offline RL methods.
To summarize, our contributions in this paper are fourfold: (i) we propose a novel generative auto-bidding method, AIGB-Pearl, that enables continuous improvement in generation quality through exploration beyond the offline dataset; (ii) we propose a provable KL-Lipschitz constrained score maximization objective with a sub-optimality bound, ensuring a safe and effective generalization for AIGB; (iii) we devise a practical algorithm with synchronous coupling that effectively ensures the Lipschitz requirement for the generative model; (iv) extensive simulated and real-world experiments demonstrate that AIGB-Pearl achieves SOTA performance and verify the effectiveness of the developed techniques in enhancing safe and effective generalization.
2 Preliminaries
2.1 Problem Statement
This work studies the auto-bidding problem for a single advertiser subject to a budget constraint. The auction mechanism follows a sealed-bid, second-price rule, and the objective is to devise a bidding policy that maximizes the cumulative value of the impressions won over a finite bidding episode, such as a day. As established in (He et al., 2021), the optimal bid for each impression $i$ is proportional to its intrinsic value $v_{i}>0$, scaled by a non-negative factor $a\geq 0$ that remains consistent across all impressions. Under this strategy, the advertiser wins an impression $i$ if $av_{i}\geq p_{i}$, where $p_{i}>0$ denotes the market price, and pays $p_{i}$ upon winning. The Return on Investment (ROI) of impression $i$ is defined as $v_{i}/p_{i}$, and we denote its upper bound as $R_{m}\triangleq\max_{i}v_{i}/p_{i}$.
However, the scaling factor is unknown in advance, and the volatility of impressions drives its continual change throughout the bidding process. Hence, a standard practice involves recalibrating the scaling factor $a$ at fixed intervals of $T\in\mathbb{N}_{+}$ time steps (Guo et al., 2024; He et al., 2021; Mou et al., 2022). This casts the auto-bidding to a sequential decision-making problem.
Specifically, the auto-bidding problem can be modeled as a Markov Decision Process (MDP) $\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P}>$. The state $s_{t}\triangleq[t,c_{t-1},x]\in\mathcal{S}$ is composed of the current time step $t\in[T]$, the cost $c_{t-1}>0$ of impressions won between time step t1t-1 and $t$, and the static individual feature $x$ of the advertiser. The action $a_{t}\in\mathcal{A}$ denotes the calibrated scaling factor at time step $t$. The reward $r_{t}>=0$ describes the value of the impressions won between time steps $t$ and $t+1$, and $\mathcal{P}$ denotes the state transition rule. The auto-bidding problem can be formulated as:
$\max_{a_{1},a_{2},\cdots,a_{T}}\;\mathbb{E}_{s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})}\bigg[\sum_{t=1}^{T}r_{t}\bigg],\;\;\mathrm{s.t.}\sum_{t=1}^{T}c_{t}\leq B,$ (1)
where $B>0$ denotes the budget of the advertiser.
Offline Setting. Due to safety concerns—common in real-world advertising systems—we are restricted to learning the optimal bidding policy from a static offline dataset $\mathcal{D}$ composed of historical states and actions along with associated rewards. This makes the considered auto-bidding problem an offline sequential decision-making task.
2.2 Offline RL Methods
RL constitutes a standard approach for auto-bidding problems, seeking an optimal bidding policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes cumulative rewards. Specifically, this is typically achieved by learning a Q-value function, $Q(s_{t},a_{t})\triangleq\mathbb{E}_{\pi}[\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}]$, through temporal difference (TD) error minimization:
$\min_{Q}\;\mathbb{E}_{(s_{t},a_{t},r_{t},s_{t+1})\sim\mathcal{D}}[Q(s_{t},a_{t})-r_{t}-\max_{a_{t+1}}\hat{Q}(s_{t+1},a_{t+1})]^{2},$ (2)
where $\hat{Q}$ is a target Q-value function with parameters updated via Polyak averaging (Mnih et al., 2015). Upon convergence, the optimal bidding policy is derived as $\pi(s_{t})=\arg\max_{a_{t}}Q(s_{t},a_{t})$.
Due to the offline setting of the considered auto-bidding problem, directly employing Eq. 2 results in the infamous out-of-distribution (OOD) problem (Fujimoto et al., 2019), making the policy erroneously deviate from the offline dataset $\mathcal{D}$. As a standard solution, offline RL (Yu et al., 2020; Kumar et al., 2020; Kidambi et al., 2020; Wang et al., 2022) constrains the policy’s behavior near $\mathcal{D}$ during TD learning, enabling reliable generalization beyond the offline dataset.
However, offline RL methods notoriously suffer from training instability caused by TD-learning (Peng et al., 2024), where the bootstrapped value of the Q function serves as its training label, resulting in an erroneous ground truth. Training stability is critical in auto-bidding due to the absence of an accurate offline policy evaluation method and the high cost of online policy examination in a real-world advertising system (Mou et al., 2022).
2.3 Generative Auto-bidding Methods
Definition 1 (Trajectory and Trajectory Quality).
The trajectory is formalized as the state sequence throughout the bidding episode, i.e., $\tau\triangleq[s_{1},s_{2},\cdots,s_{T}]$. The trajectory quality is defined as the cumulative reward of the trajectory, i.e., $y(\tau)\triangleq\sum_{t=1}^{T}r_{t}$  [^1] Note that, as in real-world advertising systems, the bidding process will automatically suspend once the advertiser’s budget runs out, and thereby, any action sequence will not violate the budget constraint. Hence, the trajectory quality can be directly defined as the cumulative reward of the trajectory..
Unlike RL methods, the AI-generated auto-bidding (AIGB) (Guo et al., 2024) treats the auto-bidding problem as a sequence generation task. Specifically, a generative model is employed to fit the conditional trajectory distribution $p_{\theta}(\tau|y(\tau))$ within the offline dataset $\mathcal{D}$, i.e.,
$\max_{\theta}\;\mathbb{E}_{(\tau,y(\tau))\sim\mathcal{D}}[\log p_{\theta}(\tau|y(\tau))],$ (3)
where $\theta$ denotes the parameter. Let $y_{m}>0$ be the maximum trajectory quality in $\mathcal{D}$, we have $\forall y\in\mathcal{D},y\in[0,y_{m}]$. During inference, AIGB follows a planning-and-control architecture. Specifically, at each time step, a trajectory is sampled from the trained generative model that acts as the planner, with a manually set condition $y^{*}\triangleq(1+\epsilon)y_{m}$, where $\epsilon>0$ is a hyper-parameter with usually a small value. Then, an extra off-the-shelf inverse dynamic model (Agrawal et al., 2016), acting as the controller, is employed to compute the action. See Appendix B for detailed descriptions. AIGB avoids TD learning and generally outperforms offline RL methods (Guo et al., 2024).
However, the modeling approach in AIGB modeling does not explicitly align with the performance optimization objective of the auto-bidding problem. As a result, AIGB primarily relies on imitating trajectories from the offline dataset, lacking the ability to actively explore higher-quality trajectories and improve its generation quality from the performance feedback. This limitation imposes a fundamental performance ceiling. Particularly, as generative models (e.g., diffusion models) are prone to overfitting with limited data (Ajay et al., 2022), their generalization ability beyond the offline dataset can be restricted, especially when there is a lack of high-quality trajectories in the offline dataset.
3 Method
Enabling AIGB to explore higher-quality trajectories actively beyond the offline dataset can enhance its performance and generalization ability. To this end, we propose AIGB-Pearl (Planning with EvaluAtor via RL) that constructs a trajectory evaluator (referred to as the evaluator for simplicity) to integrate RL methods into AIGB’s planner training. Specifically, the evaluator learns a score $\hat{y}_{\phi}(\tau)$ to estimate the trajectory quality $y(\tau)$ via supervised learning, where $\phi$ denotes the evaluator parameter. Then, with $\phi$ fixed, the planner tries to maximize the score of its generation through constant interactions with the evaluator, as shown in Fig. 1. Formally, this can be formulated as:
$\max_{\theta}\;L(\theta)\triangleq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[\hat{y}_{\phi}(\tau)],$ (4)
where the condition is fixed to $y^{*}$ in both training and inference stages to ensure consistency.
The effectiveness of AIGB-Pearl fundamentally hinges on the evaluator’s reliability. However, given the offline setting of the considered auto-bidding problem, the evaluator learning is confined to the offline dataset $\mathcal{D}$, and directly solving Eq. 4 can induce the infamous OOD problem due to the evaluator’s generalization limits, potentially resulting in the planner’s true performance degradation.
To address this challenge, we begin by examining the theoretical bounds on the evaluator’s bias. Then, guided by this analysis, we propose a KL-Lipschitz-constrained score maximization objective for the planner to ensure reliable evaluator utilization in Section 3.1. This objective is theoretically justified by a sub-optimality bound, as established in Section 3.1.1. Finally, a practical algorithm is presented in Section 3.2, where a synchronous coupling method is employed to realize the planner’s Lipschitz constraint.
3.1 KL-Lipschitz-constrained score maximization
This section focuses on the reliable exploitation of the evaluator-guided score maximization.
Our basic idea is to optimize $\theta$ within a domain where the gap between the planner’s score $L(\theta)$ and its true performance $J(\theta)\triangleq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)]$ is bounded by a small certifiable upper bound. This ensures the score maximization occurs only in regions where the evaluator is reliable.
$|J(\theta)-L(\theta)|=|\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)]-\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[\hat{y}_{\phi}(\tau)]|.$ (5)
In the following, we investigate this gap. Specifically, we find that the trajectory quality $y(\tau)$ is a Lipschitz continuous function as stated in Theorem 1, and the proof is given in Appendix C.1.
Theorem 1 (Lipschitz Continuous of $y(\tau)$.).
The trajectory quality $y(\tau)$ is $\sqrt{T}R_{m}$-Lipschitz continuous with respect to the Frobenius norm.
Based on this property, we incorporate a $\sqrt{T}R_{m}$-Lipschitz continuity constraint into the evaluator’s design (as described in Section 3.2.1), and denote the Lipschitz constant of the resulting evaluator as $k\sqrt{T}R_{m}$, where $k\geq 0$ reflects the violation degree. A tighter satisfaction of the Lipschitz constraint by the evaluator results in $k$ closer to $1$.
Equipped with Theorem 1 and the Lipschitz property of the evaluator, we derive the following upper bound on the performance gap, and the proof is given in Appendix C.2.
Theorem 2 (Evaluator Bias in Planning Performance Bound).
Let the upper bound of the evaluator’s bias on its training dataset $\mathcal{D}$ be $\delta_{D}>0$. The gap between the planner’s score $L(\theta)$ and its true performance $J(\theta)$ can be bounded by:
$|J(\theta)-L(\theta)|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}\bigg[\underbrace{W_{1}(p_{\theta}(\tau|y^{*}),p_{\theta}(\tau|y))}_{\text{ Lipschitz sensitivity to $y$}}+\underbrace{W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y))}_{\text{imitation error on $\mathcal{D}$}}\bigg],$
where $W_{1}$ denotes the 1-Wasserstein distance.
Note that $\delta_{D}$ could be regulated to a small value via supervised training of the evaluator. Consequently, bounding the planner’s performance gap requires constraining the following two factors:
  • the planner’s Lipschitz sensitivity to condition $y$ (the first Wasserstein term)
  • the planner’s imitation error on the offline dataset (the second Wasserstein term).
Formally, we establish that the expectation of the first Wasserstein term can be bounded by the Lipschitz constant $\text{Lip}_{W_{1}}(p_{\theta}(\tau|y))$ of the planner with respect to the condition $y$ measured by $W_{1}$:
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta}(\tau|y^{*}),p_{\theta}(\tau|y))]\leq(1+\epsilon)y_{m}\text{Lip}_{W_{1}}(p_{\theta}(\tau|y)).$ (6)
The proof is given in Appendix C.3. Therefore, we constrain the planner’s Lipschitz constant to a positive value $L_{p}>0$ to ensure the boundedness of the first Wasserstein term, where $L_{p}$ is a hyperparameter whose recommended range is provided later in Eq. 10.
Moreover, we establish that a constrained KL divergence $\mathbb{E}_{y\sim p_{D}(y)}[D_{\text{KL}}(p_{\theta}(\tau|y)\|p_{D}(\tau|y))]\leq\delta_{K}$ could bound the expectation of the second Wasserstein distance term as below under a mild  [^2] Note that Assumption 1 could be satisfied through strategic adjustment of the trajectory sampling probability per condition $y$ in the offline dataset $\mathcal{D}$. Here, the sampling probability adjustment should be minimal to preserve the essential characteristics of the original distribution. Formally, this can be formulated as $\min_{p_{D}^{\prime}}D_{\text{KL}}(p_{D}(\tau|y)\|p_{D}^{\prime}(\tau|y)),\;\mathrm{s.t.}\;p_{D}^{\prime}(\tau|y)\text{ satisfies LSI, }\forall y$. Assumption 1, where $\delta_{K}>0$ is a hyperparameter and can be set to a small value, close to zero. See Appendix C.4 for the proof. Note that the KL constraint inherently requires the planner to perform conditional behavior cloning on the offline dataset $\mathcal{D}$ (Guo et al., 2025).
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y))]\leq\sqrt{2\delta_{K}/\rho}.$ (7)
Assumption 1 (LSI of the Offline Dataset).
The offline data distribution $p_{D}(\tau|y)$ satisfies Logarithmic Sobolev Inequality (LSI) with constant $\rho>0$, (e.g., a Gaussian distribution where $\rho=1$).
Collectively, to effectively perform score maximization with a small and certifiable evaluator bias, we enforce Lipschitz continuity of the planner with respect to the condition $y$ while preserving its behavior cloning fidelity to the offline dataset $\mathcal{D}$. Formally, Eq. 4 is transformed to:
$\max_{\theta}\;\;\;\;L(\theta)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;\;\;\text{(Score Maximization)}$ (8)
$\;\mathrm{s.t.}\;\;\;\;\;\mathbb{E}_{y\sim p_{D}(y)}[D_{\text{KL}}(p_{\theta}(\tau|y)\|p_{D}(\tau|y))]\leq\delta_{K}\;\;\text{(KL Constraint)}$ (8a)
$\quad\;\;\;\;\;\;\;\;\text{Lip}_{W_{1}}(p_{\theta}(\tau|y))\leq L_{p}\qquad\qquad\qquad\qquad\text{(Lipschitz Constraint)}$ (8b)
Eq. 8 forms the score maximization objective in AIGB-Pearl.
Remark 1.
Intuitively, the KL and Lipschitz constraints collectively ensure the planner’s generation under condition $y^{*}$ remains within a certified neighborhood of the offline dataset $\mathcal{D}$, while the evaluator maintains high accuracy within this $\mathcal{D}$-proximal region as it is trained with the offline dataset. Therefore, Eq. 8 can realize reliable evaluator-guided score maximization.
3.1.1 sub-optimality gap bound
This section focuses on presenting and analyzing the sub-optimality bound of the solution to the proposed Eq. 8. Specifically, denote the solution to the true performance $J(\theta)$ as $\theta^{*}\triangleq\arg\max_{\theta}J(\theta)$, and denote the solution to the proposed Eq. 8 as $\hat{\theta}$. The following theorem gives the sub-optimality bound of the planner’s performance, and the proof is given in Appendix C.5.
Theorem 3 (Sub-optimality Gap Bound).
Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ be the expected distance between the optimal trajectory distribution and the trajectory distribution of the offline dataset $\mathcal{D}$. The true performance gap between the optimal parameter $\theta^{*}$ and the solution $\hat{\theta}$ to Eq. 8 is bounded by:
$J(\theta^{*})-J(\hat{\theta})\leq 2\delta_{D}+(1+2k)\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+\sqrt{2\delta_{K}/\rho}+(1+\epsilon)y_{m}L_{p}\bigg].$ (9)
Theoretical Result Analysis. In Theorem 3, the constants $\delta_{M},R_{m},T,y_{m},\epsilon,\rho$ characterize domain-specific properties of the auto-bidding task and the offline dataset $\mathcal{D}$. Nonetheless, a lower training error $\delta_{D}$ of the evaluator and a closer $k$ to $1$ correspond to a smaller sub-optimality gap. Note that $k$ cannot approach zero without compromising $\delta_{D}$, as excessively small $k$ prevents the evaluator from fitting the offline dataset $\mathcal{D}$.
Moreover, in Theorem 3, a lower behavior cloning error $\delta_{K}$ and a lower Lipschitz constant $L_{p}$ of the planner lead to a smaller sub-optimality gap. However, an excessively small $L_{p}$ prevents the planner from behavior cloning the offline dataset $\mathcal{D}$ (as the KL constraint requires), resulting in a large $\delta_{K}$. Actually, a theoretical lower bound of $L_{p}$ is the Lipschitz constant of the offline dataset distribution:
$L_{p}\geq\sup_{\begin{subarray}{c}y_{1}\neq y_{2}\end{subarray}}\frac{W_{1}(p_{D}(\tau|y_{1}),p_{D}(\tau|y_{2}))}{|y_{1}-y_{2}|}.$ (10)
3.2 Practical Algorithm Design
This section focuses on the practical algorithm implementation of Eq. 8. Section 3.2.1 first presents our reliability-enhanced evaluator architecture, followed by the synchronous-coupling-based Lipschitz planner design in Section 3.2.2.
3.2.1 Lipschitz Trajectory Evaluator
As shown in Fig. 1, the evaluator processes the trajectory $\tau$ to predict a score $\hat{y}_{\phi}(\tau)$ for quality estimation. The evaluator is trained through supervised learning, leveraging the offline dataset. Besides, to satisfy $\sqrt{T}R_{m}$-Lipschitz constraint requirement according to Theorem 2, we add Lipschitz regularization term to the training loss $l_{e}(\phi)$:
$l_{e}(\phi)=\underbrace{\mathbb{E}_{\tau\sim\mathcal{D}}\bigg[(\hat{y}_{\phi}(\tau)-y(\tau))^{2}\bigg]}_{\text{fitting the ground truth}}+\beta_{1}\underbrace{\mathbb{E}_{\tau_{1},\tau_{2}}\bigg[|\hat{y}_{\phi}(\tau_{1})-\hat{y}_{\phi}(\tau_{2})|-\sqrt{T}R_{m}\left\|\tau_{1}-\tau_{2}\right\|_{F}\bigg]_{+}}_{\text{Lipschitz Penalty}},$ (11)
where $\beta_{1}>0$ is a hyper-parameter, $\cdot]_{+}\triangleq\max\{0,\cdot\}$.
3.2.2 Lipschitz Planner With Synchronous Coupling
As shown in Fig. 1, the planner is implemented by a Causal Transformer (Chen et al., 2021b) that generates trajectories in an auto-aggressive manner. Specifically, the condition $y$ and each state in the trajectory $\tau$ are treated as input tokens, and each output state is modeled as a Gaussian distribution, $p_{\theta}(s_{t+1}|s_{1:t},y)=\mathcal{N}(\mu_{\theta}(s_{1:t},y,t),\sigma_{\theta}(s_{1:t},y,t))$. During the generation process, each output state is sampled from the Gaussian distribution using the reparameterization trick, i.e., $s_{t+1}=\mu_{\theta}(s_{1:t},y,t)+\sigma_{\theta}(s_{1:t},y,t)\cdot\eta_{t}$, where $\eta_{t}\sim\mathcal{N}(0,I)$  [^3] Note that the state $s_{t}=[t,c_{t-1},x]$ involves a time step $t$ and a static advertiser feature $x$ that do not need to be generated. Therefore, we only generate the cost $c_{t}$ in practice..
Regularized Planner Training Loss. To perform the score maximization in Eq. 8, we involve two regularization terms in the planner’s training loss $l_{p}(\theta)$, including a conditional behavior cloning loss, corresponding to the KL constraint Eq. 8a, and a Lipschitz penalty loss, corresponding to the Lipschitz constraint Eq. 8b, i.e.,
$l_{p}(\theta)=$ $-\underbrace{\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[\hat{y}_{\phi}(\tau)]}_{\text{planner score }L(\theta)}+\beta_{2}\underbrace{\mathbb{E}_{(\tau,y)\sim p_{D}}[\log p_{\theta}(\tau|y)]}_{\text{conditional behavior clone}}$
$+\beta_{3}\underbrace{\mathbb{E}_{y_{1},y_{2}\in\mathcal{D}\cup\{y^{*}\}}\bigg[{W}_{1}(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))-L_{p}|y_{1}-y_{2}|\bigg]_{+}}_{\text{Lipschitz penalty, where ${W}_{1}(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ is replaced by $\hat{W}_{1}(y_{1},y_{2};\theta)$}},$ (12)
where $\beta_{2},\beta_{3}>0$ are two hyper-parameters. With prior RL works (Sutton et al., 1999), we can derive the closed-form expression of planner’s score gradient $\nabla_{\theta}L(\theta)$ as shown in Appendix C.6. The core of the planner loss lies in the computation of ${W}_{1}(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$.
Wasserstein Upper Bound as Surrogate. Accurate computation of this Wasserstein distance term is challenging, as it requires finding the optimal coupling between $p_{\theta}(\tau|y_{1})$ and $p_{\theta}(\tau|y_{2})$ that minimizes the expected transportation cost. Nonetheless, we can choose a certain coupling $\gamma\in\Gamma(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ to obtain an upper bound of this Wasserstein term, i.e.,
${W}_{1}(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ $\triangleq\inf_{\gamma\in\Gamma(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))}\mathbb{E}_{(\tau_{1},\tau_{2})\sim\gamma}\bigg[\sum_{t}\|s_{t}^{1}-s_{t}^{2}\|\bigg]$
$\leq\mathbb{E}_{\eta_{1:T}\sim\mathcal{N}(0,I)}\bigg[\sum_{t}\|s_{t}^{1}-s_{t}^{2}]\|\bigg]\triangleq\hat{W}_{1}(y_{1},y_{2};\theta).$ (13)
where $\hat{W}_{1}(y_{1},y_{2};\theta)$ denotes the upper bound, and $s^{i}_{t}$ is the $t$-th state in trajectory $\tau_{i}$. It can be seen that $\hat{W}_{1}(y_{1},y_{2};\theta)\leq L_{p}|y_{1}-y_{2}|$ acts as a sufficient condition to make the planner $L_{p}$-Lipschitz continuous. Thus, we replace ${W}_{1}(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ by this upper bound in the planner loss.
Synchronous Coupling Wasserstein. Instead of using random couplings, we employ a synchronous coupling $\gamma_{\text{sync}}$ to make the upper bound tighter. Specifically, two trajectories $\tau_{1}$ and $\tau_{2}$—conditioned on $y_{1}$ and $y_{2}$, respectively—are generated using the same sequence of Gaussian noise $\{\eta_{1},\eta_{2},...\eta_{T}\}$. The definition of $\hat{W}_{1}(y_{1},y_{2};\theta)$ is given in Eq. 3.2.2. Compared to random couplings, the synchronous coupling reduces spurious variance in the trajectory comparison by aligning stochasticity through shared noise, resulting in a tighter upper bound on the Wasserstein distance (Lindvall, 2002).
Moreover, if we fix the predicted variance $\sigma_{\theta}$ of the planner as a fixed constant, then the expression of $\hat{W}_{1}(y_{1},y_{2};\theta)$ can be further simplified to $\hat{W}_{1}(y_{1},y_{2};\theta)=\sum_{t}\|\mu_{\theta}(s^{1}_{1:t},y_{1},t)-\mu_{\theta}(s^{2}_{1:t},y_{2},t)\|$. The overall AIGB-Pearl algorithm is summarized in Algorithm 1 in Appendix D due to page limits.
4 Experiments
We conduct both simulated and real-world experiments to validate the effectiveness of our approach. In the experiments, we mainly investigate the following Research Questions (RQs): (1) Does enhancing AIGB with policy optimization improve overall performance, and can it generalize better to unseen data compared to existing AIGB methods? (Section 4.2) (2) How does the KL-Lipschitz constraint affect the performance of the planner? (Section 4.3) (3) Can the proposed method guarantee the Lipschitz property of the evaluator and the planner? (Section 4.4). The training stability of AIGB-Pearl is studied in Appendix E.5.
Figure 1: AIGB-Pearl (Planning with EvaluAtor via RL) constructs a trajectory evaluator to score the trajectory quality and let the planner maximize the obtained score under the KL-Lipschitz constraint through continuous interaction with the evaluator. A synchronous coupling method is used to estimate the Wasserstein term in the Lipschitz penalty.
4.1 Experiment Setup
Experiment Environment. We conduct simulated experiments in an open-source offline advertising system with 30 advertisers of four budget levels (1.5k, 2.0k, 2.5k, and 3.0k), as in (Mou et al., 2022; Guo et al., 2024). The offline dataset comprises 5k trajectories generated by 20 advertisers. Extra detailed settings of simulated experiments are given in Appendix E.1. For real-world experiments, we conduct online A/B tests on one of the world’s largest E-commerce platforms, TaoBao. The offline dataset comprises 200k trajectories of 10k advertisers. See Appendix E.2 for extra detailed settings of real-world experiments. In both simulated and real-world experiments, we employ the same inverse dynamic model (Agrawal et al., 2016) used in the AIGB as the controller.
Baselines. We compare our method with the state-of-the-art AIGB methods, including DiffBid (Guo et al., 2024) and DT (Chen et al., 2021a) that learn from conditional behavior cloning of the offline dataset with a diffusion model and a Causal Transformer, respectively. We also compare our method with RL auto-bidding methods, including USCB (He et al., 2021) that learns the auto-bidding policy in a manually constructed advertising system with DDPG (Silver et al., 2014); and offline RL auto-bidding methods, including model-free offline RL methods BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022) and Diff-QL (Wang et al., 2022), and model-based offline RL method MOPO (Yu et al., 2020).
Performance Index. The objective in the auto-bidding problem Eq. 1, i.e., the cumulative rewards over the bidding episode, acts as the main performance index in our experiments and is referred to as the gross merchandise volume, GMV. In addition, we utilize three other metrics commonly used in the auto-bidding problem to evaluate the performance of our approach. The first metric is the total number of impressions won over the bidding episode, referred to as the BuyCnt. The second metric is the Cost over the bidding episode, and the third one is the return on investment ROI defined as the ratio between the GMV and the Cost. Note that larger values of GMV, BuyCnt, and ROI with a Cost oscillating within an acceptable tolerance ($\pm 2\%$) indicate a better performance.
4.2 Overall Performance
Table 1: Overall performance (GMV) in simulated experiments with 30 advertisers. $\Delta$ indicates the relative improvement of AIGB-Pearl against the most competitive baseline (which is underlined). Note that the absolute values are normalized without specific meanings; only $\Delta$ matters.
Budget USCB BCQ CQL IQL Diff-QL MOPO DT DiffBid AIGB-Pearl $\bm{\Delta}$
1.5k 454.25 454.72 461.82 456.80 469.73 470.38 477.39 480.76 502.98 +4.62%
2.0k 482.67 483.50 475.78 486.56 487.91 489.27 507.30 511.17 521.84 +2.09%
2.5k 497.66 498.77 481.37 518.27 510.83 523.91 527.88 531.29 545.03 +2.59%
3.0k 500.60 501.86 491.36 549.19 552.73 549.01 550.66 556.32 574.17 +3.21%
Table 2: Overall performance in real-world A/B tests, involving 6k advertisers over 19 days.
Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DiffBid 76,390,174 650,962 5.31 14,395,290 USCB 52,182,805 516,994 4.92 10,598,486 AIGB-Pearl 78,676,009 665,173 5.41 14,551,054 AIGB-Pearl 53,973,101 520,796 5.13 10,515,772 $\bm{\Delta}$ +3.00% +2.20% +1.89% +1.10% $\bm{\Delta}$ +3.43% +0.74% +4.24% -0.78% Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DT 34,808,665 341,995 5.61 6,205,665 MOPO 51,674,071 579,332 3.08 16,771,892 AIGB-Pearl 35,957,933 344,194 5.77 6,246,512 AIGB-Pearl 53,292,945 591,741 3.23 16,475,670 $\bm{\Delta}$ +3.30% +0.64% +0.16% +0.66% $\bm{\Delta}$ +3.13% +2.14% +4.87% -1.77%
Table 3: Generalization performance in real-world A/B tests with unseen advertisers against AIGB methods, involving 4k advertisers over 19 days.
Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DiffBid 67,092,973 553,020 5.39 12,444,306 DT 30,562,007 300,271 5.61 5,450,573 AIGB-Pearl 69,252,539 565,776 5.53 12,534,379 AIGB-Pearl 31,502,309 305,202 5.74 5,484,473 $\bm{\Delta}$ +3.32% +2.31% +2.48% +0.72% $\bm{\Delta}$ +3.08% +1.64% +2.32% +0.62%
To answer RQ(1): Table 1 shows that our method consistently outperforms all baselines in GMV across all four budget levels in simulated experiments. In real-world experiments, the results in Table 2 show that our method also achieves superior performance in GMV, BuyCnt, and ROI, with Cost fluctuations within 2%. Notably, both simulated and real-world experiments consistently demonstrate that AIGB-Pearl achieves a +3% improvement in GMV over the AIGB, the state-of-the-art auto-bidding method. Since our method and DiffBid share the same controller, the performance gain stems solely from the planner. This provides strong empirical evidence that the proposed conservative RL learning for score maximization effectively enhances overall performance.
Notably, we also apply AIGB-Pearl to another important auto-bidding problem, named TargetROAS. As shown in Appendix E.3, real-world experiments show that AIGB-Pearl achieves a +5% improvement in GMV compared to AIGB.
Generalization Ability. We examine the performance of AIGB-Pearl on advertisers that are not used to generate trajectories in the offline dataset, comparing it against existing AIGB methods. For simplicity, we refer to these advertisers as advertisers outside the offline dataset. Table 3 reports the performance on 4k advertisers outside the offline dataset in real-world experiments. AIGB-Pearl demonstrates consistently better results in terms of GMV (+3%), BuyCnt, and ROI, while maintaining Cost fluctuations within 2% compared to baselines. These results indicate that the proposed method has better generalization ability than existing AIGB methods.
4.3 Ablation Study
To answer RQ(2): We remove the KL constraint and the Lipschitz constraint from AIGB-Pearl individually and evaluate the model’s performance in each ablated variant with real-world A/B tests. The results are presented in Table 4. It can be seen that the KL constraint contributes +1.1% improvement in GMV, and the Lipschitz constraint provides +1.8% improvement in GMV, demonstrating their respective roles in enhancing AIGB-Pearl’s performance.
Table 4: Ablation Study. The effectiveness of the KL constraint and the Lipschitz constraint in Real-world A/B tests, involving 6k advertisers over 8 days.
AIGB-Pearl GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost w/o KL 30,906,963 292,605 4.25 7,269,018 w/o Lipschitz 32,284,972 268,551 5.73 5,634,304 with KL 31,243,688 292,783 4.26 7,342,485 with Lipschitz 32,869,329 281,979 5.79 5,678,252 $\bm{\Delta}$ +1.09% +0.06% +0.08% +1.01% $\bm{\Delta}$ +1.81% +0.50% +1.05% +0.78%
Visualization. Three AIGB-Pearl-generated trajectory examples are presented in Fig. 2. As can be observed, the trajectories generated by AIGB-Pearl are plausible. In contrast, the ablated variant without the KL and Lipschitz constraints produces trajectories that significantly deviate from the optimal trajectory in the offline dataset  [^4] The optimal trajectory in the offline dataset can act as a reference trajectory, and empirically, optimal trajectories should not deviate from this reference largely. and exhibit clear pathological behaviors—such as excessive budget consumption, backward-trending pacing, and underutilization of available budgets (see Appendix E.4 for explanation)—which further validate the KL-Lipschitz constraint necessity.
Figure 2: Trajectory Generation Visualization. Three cases are presented. Here, the AIGB-Pearl generates plausible trajectories, whereas its variant without the KL-Lipschitz constraint produces generations that significantly deviate from the reference and exhibit evident issues.
4.4 Lipschitz Value Examination
Figure 3: Examination of Evaluator Lipschitz.
Figure 4: Examination of Planner Lipschitz.
To answer RQ(3): We report that the Lipschitz value of the trajectory quality $y(\tau)$ and the trajectory conditional distribution $p_{D}$ of the offline dataset are $1.62$ and $38$, respectively, and we set $L_{p}=50$. To calculate the Lipschitz values of the evaluator and the planner, we sample $8,000$ pairs of trajectories and calculate their Lipschitz value. The results are shown in Fig. 4 and Fig. 4. It can be observed that most sample values satisfy the Lipschitz constraint, and the Lipschitz constants of the evaluator $\hat{y}_{\phi}(\tau)$ and planner $p_{\theta}(\tau|y)$ are $2.2$ and $56$, respectively, near $1.62$ and $50$. This indicates that the Lipschitz constraints of the evaluator and the planner are successfully satisfied.
5 Conclusions
This paper proposes AIGB-Pearl to enhance AIGB with reward evaluation and policy optimization. By introducing a trajectory evaluator and a provable KL-Lipschitz-constrained score maximization objective, our approach ensures safe and efficient generalization beyond the offline dataset, supported by theoretical guarantees. Extensive simulated and real-world experiments validate the SOTA performance of our approach.
Appendix A Related Works
A.1 RL-based Auto-bidding Methods
Auto-bidding plays a critical role in online advertising by automatically placing bids, allowing advertisers to participate efficiently in real-time auctions (Balseiro et al., 2021a; Deng et al., 2021; Balseiro et al., 2021b). The auto-bidding problem can be modeled as a Markov Decision Process and addressed using reinforcement learning techniques. USCB (He et al., 2021) proposes a unified solution to the constrained bidding problem, employing an RL method, DDPG (Silver et al., 2014), to dynamically adjust parameters to an optimal bidding strategy. Mou et al. (2022) design a sustainable online reinforcement learning framework that iteratively alternates between online explorations and offline training, alleviating the sim2rel problem. A few studies explore multi-agent RL for auto-bidding (Jin et al., 2018; Guan et al., 2021; Wen et al., 2022), while several focus on budget allocation and bidding strategies in multi-channel scenarios using RL-based approaches (Wang et al., 2023; Deng et al., 2023; Duan et al., 2025). Importantly, offline RL methods such as BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022), and MOPO (Yu et al., 2020) have demonstrated significant potential in this domain. These methods allow policy learning from pre-collected datasets without requiring online interaction. Moreover, offline RL, such as Diffusion-QL (Wang et al., 2022), adopts generative models as the policy model architecture for better expressive capacity.
However, RL-based methods often suffer from training instability caused by bootstrapping and alternating training paradigms between critics and actors. Training instability typically deteriorates policy performance (Sutton et al., 1998). Moreover, training stability is even more critical in auto-bidding considering two domain-specific challenges: the absence of an accurate offline policy evaluation method and the high cost of online policy examination in a real-world advertising system (Mou et al., 2022). Therefore, stable convergence to a well-performed policy is essential to ensure deployment reliability and system safety.
A.2 Generative Auto-bidding Methods
Generative models exhibit strong capabilities in capturing and replicating the underlying data distributions across a wide range of fields (Kingma & Welling, 2022; Goodfellow et al., 2020; Pan et al., 2023; Sohl-Dickstein et al., 2015; Ho et al., 2020; Vaswani et al., 2017). They can be effectively incorporated into decision-making systems by generating complete trajectories to guide agents toward high reward behaviors (Zhu et al., 2023; Kang et al., 2023; Li et al., 2025). In particular, Decision Transformer (DT) (Chen et al., 2021a) reframes RL as a conditional sequence modeling problem and leverages transformer architectures to generate actions conditioned on desired returns, historical states, and actions. AIGB (Guo et al., 2024) extends the generative perspective to the auto-bidding domain by formulating auto-bidding as a conditional generative modeling problem. DiffBid generates a state trajectory based on the desired return utilizing a conditional diffusion model, and then generates actions aligned with the optimized trajectory. These methods achieve superior performance in auto-bidding and offer distinct advantages over traditional RL methods. They do not rely on the bootstrapping mechanism commonly used in RL, thereby avoiding the instability caused by the deadly triad. Even so, these generative auto-bidding methods still encounter a performance bottleneck due to their neglect of fine-grained generation quality evaluation and inability to explore beyond static datasets. In contrast, our method facilitates both reward evaluation and policy search through a learned trajectory evaluator.
Appendix B AIGB Method Details
AIGB models the sequential decision-making problem through conditional diffusion modeling, enabling effective trajectory generation for auto-bidding scenarios. Specifically, AIGB utilizes the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) for generation. The forward and reverse processes are modeled as:
$q(\tau_{k+1}|\tau_{k}),\quad p_{\theta}(\tau_{k}|\tau_{k+1},y(\tau)),$ (14)
respectively, where $q$ represents the forward noising process while $p_{\theta}$ the reverse denoising process.
Forward Process. In the forward process, the noise is gradually added to the latent variable by a Markov chain with pre-defined variance schedule $\beta_{k}$:
$q(\tau_{k}|\tau_{k-1})=\mathcal{N}(\tau_{k};\sqrt{1-\beta_{k}}\tau_{k-1},\beta_{k}I)$ (15)
where $k\in[K]$ refers to the diffusion step, $\tau_{k}\triangleq[s_{1},s_{2},\cdots,{s}_{T}]_{k}$ represents the latent variable in the $k$-th diffusion step, and $\tau_{0}$ is the original trajectory. A notable property of the forward process is that $\tau_{k}$ at an arbitrary time-step $k$ can be sampled in closed form as:
$q(\tau_{k}|\tau_{0})=\mathcal{N}(\tau_{k};\sqrt{\bar{\alpha}_{k}}\tau_{0},(1-\bar{\alpha}_{k})I),$ (16)
where $\alpha_{k}=1-\beta_{k}$ and $\bar{\alpha}_{k}=\prod^{k}_{i=1}\alpha_{k}$. When $k\rightarrow\infty$, $\tau_{k}$ approaches a standard Gaussian distribution. In particular, AIGB employs a cosine noise schedule (Nichol & Dhariwal, 2021) to control the schedule $\beta_{k}$.
Reverse Process. In the reverse process, diffusion models aim to remove the added noise on $\tau_{K}$ and recursively recover $\tau_{0}$. This process is governed by the conditional model $p_{\theta}(\tau_{k-1}|\tau_{k},y(\tau))$, which is parameterized through a noise prediction model $\epsilon_{\theta}(\tau_{k},y(\tau),k)$. AIGB adopts a temporal U-Net (Ronneberger et al., 2015) for the noise prediction model, a common choice in diffusion-based decision-making methods (Ajay et al., 2022).
B.1 Training Stage
The training of the diffusion model is typically formulated as minimizing the mean squared error between the predicted noise $\epsilon_{\theta}$ and the true noise applied during the forward diffusion process. Specifically, during each iteration, we randomly sample a trajectory from the offline dataset $\mathcal{D}$ and pick a time step $t\in[T]$. We recursively add the Gaussian noise $\epsilon$ to the states in $\tau$ with time steps bigger than $t$ and predict the added noises with $\epsilon_{\theta}(\tau_{k},y(\tau),k)$, where the states between 0 and $t$ in $\tau_{k}$ are set to real history states $s_{1},s_{2},\cdots,s_{t}$. In addition to this standard objective, AIGB also incorporates a supervised loss that measures the discrepancy between the true actions and the actions predicted by an inverse dynamics model $\hat{f}_{\phi}(s_{t},\hat{s}_{t+1})$. Overall, the complete training objective of AIGB can be expressed as:
[…]
During training, the condition $y(\tau)$ is randomly dropped to enhance model robustness. This technique ensures that both the unconditional model $\epsilon_{\theta}(\tau_{k},k)$ and the conditional model $\epsilon_{\theta}(\tau_{k},y(\tau),k)$ are effectively trained together.
B.2 Inference Stage
Starting with Gaussian noise, trajectories are iteratively generated through a series of denoising steps. Specifically, AIGB uses a classifier-free guidance strategy (Ho & Salimans, 2021) to guide the generation of bidding and extract high-likelihood trajectories in the dataset. During generation, AIGB combines conditional and unconditional score estimates linearly:
$\hat{\epsilon}_{k}:=\epsilon_{\theta}(\tau_{k},k)+\omega\left(\epsilon_{\theta}\left(\tau_{k},y(\tau),k\right)-\epsilon_{\theta}\left(\tau_{k},k\right)\right),$ (18)
where $\omega$ is the guidance scale that controls the influence of the condition $y(\tau)$. This formulation effectively steers the trajectory generation towards regions of the data distribution most consistent with the given condition. The predicted state at each step is sampled from $p_{\theta}(\tau_{k-1}|\tau_{k},y(\tau))$:
$\tau_{k-1}\sim\mathcal{N}\left(\tau_{k-1}|{\mu}_{\theta}\left(\tau_{k},y(\tau),k\right),{\Sigma}_{\theta}\left(\tau_{k},k\right)\right),$ (19)
with mean and variance defined as ${\mu}_{\theta}(\tau_{k},y(\tau),k)=\frac{1}{\sqrt{\alpha_{k}}}(\tau_{k}-\frac{\beta_{k}}{\sqrt{1-\overline{\alpha}_{k}}}\hat{\epsilon}_{k})$ and $\Sigma_{\theta}(\cdot)=\beta_{k}$. Note that the initial noisy trajectory $\tau^{\prime}_{K}\sim\mathcal{N}(0,I)$ is assigned with history states $s_{1:t}$ for the first $t$ states to ensure history consistency. This is consistent with the training process. By recursively applying the reverse diffusion process using:
$\tau^{\prime}_{k-1}={\mu}_{\theta}(\tau^{\prime}_{k},y(\tau),k)+\sqrt{\beta_{k}}{z},$ (20)
where $z\sim\mathcal{N}(0,I)$, we obtain the final denoised trajectory $\tau^{\prime}_{0}$, from which the next state $\hat{s}_{t+1}$ is derived. Then the action is generated through an inverse dynamics $\hat{a}_{t}=\hat{f}_{\phi}(s_{t},\hat{s}_{t+1})$.
Appendix C Theoretical Proofs
C.1 Proof of Theorem 1.
Figure 5: The impression opportunities within time step $t$ and $t+1$, where $p_{t}^{i}/v_{t}^{i}$ is the $1/\text{ROI}$ of impression ii. Without loss of generality, consider two actions $a_{1,t}$ and $a_{2,t}$, and let $a_{2,t}\geq a_{1,t}$. The impressions within the shadow area are the impressions won by action $a_{2,t}$ but lost by action $a_{1,t}$.
Theorem 1 (Lipschitz Continuous of $y(\tau)$.).
The trajectory quality $y(\tau)$ is $\sqrt{T}R_{m}$-Lipschitz continuous with respect to the Frobenius norm.
Proof.
Recall from Section 2.1 that the cost $c_{t}$ and reward $r_{t}$ under action $a_{t}$ between time step $t$ and $t+1$ can be written as:
$c_{t}=\sum_{i}\mathbbm{1}\bigg\{a_{t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\bigg\}p_{t}^{i}\quad\text{and}\quad r_{t}=\sum_{i}\mathbbm{1}\bigg\{a_{t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\bigg\}v_{t}^{i},$ (21)
where $p_{t}^{i}$ and $v_{t}^{i}$ denote the market price and the value of the $i$-th impression between time step $t$ and $t+1$. Consider two trajectories $\tau_{1}$ and $\tau_{2}$ with action, cost and reward sequences $\{a_{1,t},c_{1,t},r_{1,t}\}_{t=1}^{T}$ and $\{a_{2,t},c_{2,t},r_{2,t}\}_{t=1}^{T}$, respectively. The trajectory quality gap between $\tau_{1}$ and $\tau_{2}$ holds that:
$|y(\tau_{1})-y(\tau_{2})|=\big|\sum_{t}r_{1,t}-\sum_{t}r_{2,t}\big|\leq\sum_{t}|r_{1,t}-r_{2,t}|.$ (22)
Consider the reward gap between time step $t$ and $t+1$, as shown in Fig.5. Without loss of generality, let $a_{2,t}\geq a_{1,t}$. We have:
$|r_{1,t}-r_{2,t}|$ $=\sum_{i}\bigg[\mathbbm{1}\bigg\{a_{2,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\bigg\}-\mathbbm{1}\bigg\{a_{1,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\bigg\}\bigg]v_{t}^{i}$
$=\sum_{i}\mathbbm{1}\bigg\{a_{2,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\geq a_{1,t}\bigg\}v_{t}^{i}$
$=\sum_{i}\mathbbm{1}\bigg\{a_{2,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\geq a_{1,t}\bigg\}\frac{v_{t}^{i}}{p_{t}^{i}}p_{t}^{i}$
$\leq R_{m}\sum_{i}\mathbbm{1}\bigg\{a_{2,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\geq a_{1,t}\bigg\}p_{t}^{i}.$ (23)
Note that the cost gap between time step $t$ and $t+1$ can be written as:
$|c_{1,t}-c_{2,t}|=\sum_{i}\bigg[\mathbbm{1}\bigg\{a_{2,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\bigg\}-\mathbbm{1}\bigg\{a_{1,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\bigg\}\bigg]p_{t}^{i}=\sum_{i}\mathbbm{1}\bigg\{a_{2,t}\geq\frac{p_{t}^{i}}{v_{t}^{i}}\geq a_{1,t}\bigg\}p_{t}^{i}.$ (24)
Therefore, we have:
$|r_{1,t}-r_{2,t}|\leq R_{m}|c_{1,t}-c_{2,t}|.$ (25)
We examine the Frobenius norm of the gap between $\tau_{1}$ and $\tau_{2}$:
$\|\tau_{1}-\tau_{2}\|_{F}$ $=\bigg\|\begin{bmatrix}1&c_{1,0}&x\\ 2&c_{1,1}&x\\ \vdots&\vdots&\vdots\\ T&c_{1,T-1}&x\end{bmatrix}-\begin{bmatrix}1&c_{2,0}&x\\ 2&c_{2,1}&x\\ \vdots&\vdots&\vdots\\ T&c_{2,T-1}&x\end{bmatrix}\bigg\|_{F}$
$=\sqrt{\sum_{t}(c_{1,t}-c_{2,t})^{2}}$
$\geq\frac{1}{\sqrt{T}}\sum_{t}|c_{1,t}-c_{2,t}|\qquad\qquad\text{(Cauchy-Schwarz Inequality)}$ (26)
Combining Eq. 22, Eq. 25 and Eq. C.1, we can obtain that:
$|y(\tau_{1})-y(\tau_{2})|$ $\leq\sum_{t}|r_{1,t}-r_{2,t}|$
$\leq R_{m}\sum_{t}|c_{1,t}-c_{2,t}|$
$\leq\sqrt{T}R_{m}\frac{1}{\sqrt{T}}\sum_{t}|c_{1,t}-c_{2,t}|$
$\leq\sqrt{T}R_{m}\left\|\tau_{1}-\tau_{2}\right\|_{F}.$ (27)
This concludes the proof. ∎
C.2 Proof of Theorem 2
Here, we list two lemmas used in the proof of Theorem 2.
Lemma 1 (Additivity of the Lipschitz).
Let $f_{1}(x)$ and $f_{2}(x)$ be two Lipschitz continuous functions with Lipschitz constants $L_{1}>0$ and $L_{2}>0$, respectively. Then $|f_{1}(x)+f_{2}(x)|$ is also a Lipschitz continuous function, with Lipschitz constant at most $L_{1}+L_{2}$.
Proof.
Recall the Reverse Triangle Inequality states that $\forall a,b$, we have $||a|-|b||\leq|a-b|$. Then, $\forall x,y$, we have:
$||f_{1}(x)+f_{2}(x)|-|f_{1}(y)+f_{2}(y)||$ $\leq|f_{1}(x)+f_{2}(x)-f_{1}(y)-f_{2}(y)|$
$\leq|f_{1}(x)-f_{1}(y)|+|f_{2}(x)-f_{2}(y)|$
$\leq(L_{1}+L_{2})|x-y|.$ (28)
This concludes the proof. ∎
Lemma 2 (Kantorovich-Rubinstein Duality Theorem (Villani, 2021)).
Let $(X,d)$ be a metric space, and let $p$ and $q$ be two probability distributions on $X$. Let $f:X\rightarrow\mathbb{R}$ be an $L$-Lipschitz function, and $W_{1}(p,q)$ denotes the 1-Wasserstein distance between $p$ and $q$. Then we have:
$|\mathbb{E}_{x\sim p}f(x)-\mathbb{E}_{x\sim q}f(x)|\leq L\cdot W_{1}(p,q).$ (29)
We next give the proof of Theorem 2.
Theorem 2 (Evaluator Bias in Planning Performance Bound).
Let the upper bound of the evaluator’s bias on its training dataset $\mathcal{D}$ be $\delta_{D}>0$. The gap between the planner’s score $L(\theta)$ and its true performance $J(\theta)$ can be bounded by:
$|J(\theta)-L(\theta)|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}\bigg[\underbrace{W_{1}(p_{\theta}(\tau|y^{*}),p_{\theta}(\tau|y))}_{\text{ Lipschitz sensitivity to $y$}}+\underbrace{W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y))}_{\text{imitation error on $\mathcal{D}$}}\bigg],$
where $W_{1}$ denotes the 1-Wasserstein distance.
Proof.
The evaluator bias in the planner’s performance can be written as:
$|J(\theta)-L(\theta)|=|\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)-\hat{y}_{\phi}(\tau)]|\leq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}\underbrace{|y(\tau)-\hat{y}_{\phi}(\tau)|}_{\triangleq f(\tau)}$ (30)
Let $f(\tau)\triangleq|y(\tau)-\hat{y}_{\phi}(\tau)|$ be the evaluator bias in trajectory $\tau$. From Theorem 1 and Lemma 1, we know that $f(\tau)$ is a $(1+k)\sqrt{T}R_{m}$-Lipschitz continuous function. Then, we have:
$|J(\theta)-L(\theta)|$ $\leq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}f(\tau)$
$=\mathbb{E}_{y\sim p_{D}(y)}\bigg[\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}f(\tau)-\mathbb{E}_{\tau\sim p_{D}(\tau|y)}f(\tau)+\mathbb{E}_{\tau\sim p_{D}(\tau|y)}f(\tau)\bigg]$
[…]
$\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta}(\tau|y^{*}),p_{D}(\tau|y))].$ (31)
To further enhance physical intuition and simplify constraint computation, we decompose the Wasserstein distance term $W_{1}(p_{\theta}(\tau|y^{*}),p_{D}(\tau|y))$ via the triangle inequality:
$W_{1}(p_{\theta}(\tau|y^{*}),p_{D}(\tau|y))\leq W_{1}(p_{\theta}(\tau|y^{*}),p_{\theta}(\tau|y))+W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y)).$ (32)
Therefore, we have:
$|J(\theta)-L(\theta)|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}\bigg[W_{1}(p_{\theta}(\tau|y^{*}),p_{\theta}(\tau|y))+W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y))\bigg],$ (33)
This concludes the proof. ∎
C.3 Proof of Eq. 6
We give the proof of Eq. 6 as follows. Denote $\text{Lip}_{W_{1}}(p_{\theta}(\tau|y))$ as the planner’s Lipschitz constant with respect to $y$ regarding the Wasserstein distance $W_{1}$, we have:
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta}(\tau|y^{*}),p_{\theta}(\tau|y))]$ $\leq\text{Lip}_{W_{1}}(p_{\theta}(\tau|y))\mathbb{E}_{y\sim p_{D}(y)}[((1+\epsilon)y_{m}-y)]$
$=\text{Lip}_{W_{1}}(p_{\theta}(\tau|y))\int_{0}^{y_{m}}p_{D}(y)[(1+\epsilon)y_{m}-y]\mathrm{d}y$
$\leq\text{Lip}_{W_{1}}(p_{\theta}(\tau|y))\int_{0}^{y_{m}}p_{D}(y)[(1+\epsilon)y_{m}]\mathrm{d}y$
$=(1+\epsilon)y_{m}\text{Lip}_{W_{1}}(p_{\theta}(\tau|y)),$ (34)
where we leverage the non-negativity property of the condition $y\geq 0,\forall y\in\mathcal{D}$. This completes the proof.
C.4 Proof of Eq. 7
Lemma 3 (Talagrand’s inequality (Otto & Villani, 2000)).
Suppose that a probability distribution $Q$ satisfies a logarithmic Sobolev inequality (LSI) with constant $\rho>0$. Then for any probability distribution $P$ that is absolutely continuous with respect to $Q$, the following holds:
$W_{2}(P,Q)\leq\sqrt{\frac{2}{\rho}D_{KL}(P||Q)}.$ (35)
Equipped with the above lemma, as well as the KL-constraint
$\mathbb{E}_{y\sim p_{D}(y)}[D_{\text{KL}}(p_{\theta}(\tau|y)\|p_{D}(\tau|y))]\leq\delta_{K},$ (36)
we give the proof of Eq. 7 as follows. Specifically, with Assumption 1, we have:
$W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y))$ $\leq W_{2}(p_{\theta}(\tau|y),p_{D}(\tau|y))\qquad\qquad\text{(holder inequality)}$
[…]
Then, taking the expectation over $y\sim p_{D}(y)$ on both sides, we have:
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta}(\tau|y),p_{D}(\tau|y))]$ $\leq\mathbb{E}_{y\sim p_{D}(y)}\bigg[\sqrt{\frac{2}{\rho}D_{KL}(p_{\theta}(\tau|y)\|p_{D}(\tau|y))}\bigg]$
$\leq\sqrt{\frac{2}{\rho}\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta}(\tau|y)\|p_{D}(\tau|y))]}\quad\text{(Jensen Inequality)}$
[…]
This completes the proof.
C.5 Proof of Theorem 3
Theorem 3 (Sub-optimality Gap Bound).
The true performance gap between the optimal parameter $\theta^{*}$ and the solution $\hat{\theta}$ to Eq. 8 is bounded by:
$J(\theta^{*})-J(\hat{\theta})\leq 2\delta_{D}+(1+2k)\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+\sqrt{2\delta_{K}/\rho}+(1+\epsilon)y_{m}L_{p}\bigg],$ (39)
where $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ denotes the distance between the optimal trajectory distribution and the trajectory distribution of the offline dataset $\mathcal{D}$.
Proof.
The sub-optimality gap can be expressed as follows:
[…]
$\leq\underbrace{|J(\theta^{*})-L(\theta^{*})|}_{\text{evaluator bias in }p_{\theta^{*}}}+\underbrace{|L(\theta^{*})-L(\hat{\theta})|}_{\text{score gap}}+\underbrace{|L(\hat{\theta})-J(\hat{\theta})|}_{\text{evaluator bias in }p_{\hat{\theta}}}$. (40)
We examine the above three terms accordingly.
(1) Evaluator Bias in $p_{\theta^{*}}$. Denote the evaluator bias on trajectory $\tau$ as $f(\tau)\triangleq|y(\tau)-\hat{y}_{\phi}(\tau)|$. Following the derivation process in Eq. C.2, we have:
$|J(\theta^{*})-L(\theta^{*})|$ $\leq\mathbb{E}_{y\sim p_{D}(y)}\mathbb{E}_{\tau\sim p_{D}(\tau|y)}f(\tau)+\mathbb{E}_{y\sim p_{D}(y)}\bigg[\mathbb{E}_{\tau\sim p_{\theta^{*}}(\tau|y^{*})}f(\tau)-\mathbb{E}_{\tau\sim p_{D}(\tau|y)}f(\tau)\bigg]$
$\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))],$ (41)
where $W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))$ denotes the probability distribution distance between the optimal planner and the offline dataset. Based on Assumption 1 and the derivation in Eq. C.4, we have:
$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))]\leq\sqrt{\frac{2}{\rho}\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}(\tau|y)}\|p_{D}(\tau|y))]}$ (42)
Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ be the distance between the optimal trajectory distribution and the offline dataset trajectory distribution. We have:
$|J(\theta^{*})-L(\theta^{*})|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\sqrt{2\delta_{M}/\rho}.$ (43)
(2) Score Gap. Recall that $\hat{y}_{\phi}(\tau)$ is a $k\sqrt{T}R_{m}$-Lipschitz continuous function with the Lipschitz constraint design. With Lemma 2, we have:
[…]
$\leq k\sqrt{T}R_{m}W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{\hat{\theta}}(\tau|y^{*}))$
$\leq k\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}\bigg[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))+W_{1}(p_{D}(\tau|y),p_{\hat{\theta}}(\tau|y^{*}))\bigg]$
$\leq k\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg],$ (44)
where we leverage the fact that $\hat{\theta}$ satisfies the KL and Lipschitz constraint in Eq. 8, and we leverage the results in Eq. 6 and Eq. 7.
(3) Evaluator Bias in $p_{\hat{\theta}}$. Since $\hat{\theta}$ satisfies the KL and Lipschitz constraint in Eq. 8, we can use the results in Theorem 2, Eq. 6 and Eq. 7 to obtain:
$|J(\hat{\theta})-L(\hat{\theta})|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}],$ (45)
Overall, combining the results in Eq. 43, Eq. C.5, and Eq. 45, we have:
$J(\theta^{*})-J(\hat{\theta})\leq 2\delta_{D}+(1+2k)\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+\sqrt{2\delta_{K}/\rho}+(1+\epsilon)y_{m}L_{p}\bigg].$ (46)
This concludes the proof.
C.6 Proof of Score Gradient
The probability of the trajectory generated by the Causal Transformer can be decomposed into:
$p_{\theta}(s_{1:T}|y)=\Pi_{t}p_{\theta}(s_{t}|s_{1:t-1},y).$ (47)
Then, we have:
$\nabla_{\theta}L(\theta)$ $=\nabla_{\theta}\int_{\tau}p_{\theta}(\tau|y^{*})\hat{y}_{\phi}(\tau)\mathrm{d}\tau$
$=\nabla_{\theta}\int_{s_{1},\cdots,s_{T}}p_{\theta}(s_{1:T},\tau|y^{*})\hat{y}_{\phi}(\tau)\mathrm{d}s_{1}\cdots\mathrm{d}s_{T}$
$=\int_{s_{1:T}}p_{\theta}(s_{1:T}|y^{*})\frac{\nabla_{\theta}p_{\theta}(s_{1:T}|y^{*})}{p_{\theta}(s_{1:T}|y^{*})}\hat{y}_{\phi}(\tau)\mathrm{d}s_{1}\cdots\mathrm{d}s_{T}$
$=\mathbb{E}_{s_{1:T}\sim p_{\theta}(s_{1:T}|y^{*})}\bigg[\nabla_{\theta}\log p_{\theta}(s_{1:T}|y^{*})\hat{y}_{\phi}(\tau)\bigg]$
$=\mathbb{E}_{s_{1:T}\sim p_{\theta}(s_{1:T}|y^{*})}\bigg[\nabla_{\theta}\log\Pi_{t}p_{\theta}(s_{t}|s_{1:t-1},y^{*})\hat{y}_{\phi}(\tau)\bigg]$
$=\mathbb{E}_{s_{1:T}\sim p_{\theta}(s_{1:T}|y^{*})}\bigg[\sum_{t}\nabla_{\theta}\log p_{\theta}(s_{t}|s_{1:t-1},y^{*})\hat{y}_{\phi}(\tau)\bigg].$ (48)
Appendix D AIGB-Pearl Algorithm Summary
Algorithm 1 summarizes the training process of AIGB-Pearl. The development of AIGB-Pearl is supported by ROLL (Wang et al., 2025).
Input : Offline dataset $\mathcal{D}$, desired condition $y^{*}$, hyper-parameters $\beta_{1}$, $\beta_{2},\beta_{3}$.
Output : Optimized $\theta$} and $\phi$
Initialization : randomly initialized planner parameter $\theta$}, trajectory evaluator parameters $\phi$}
// Determining the Lipschitz Value
Calculate the Lipschitz value of $y(\tau)$ and $p_{D}(\tau|y)$ using the offline dataset $\mathcal{D}$.
Set the Lipschitz constraint value $L_{e}$ for the evaluator and $L_{p}$ for the planner to be bigger than the Lipschitz value of $y(\tau)$ and $p_{D}(\tau|y)$, respectively.
// Training the trajectory evaluator
while not converged do
  Update $\phi$ by minimizing Eq. 11;
end while
// Training the generative planner
Warm start with pretrained planner $p_{\theta}$;
while not converged do
   Generate bidding trajectories $\tau\sim p_{\theta}(\tau|y^{*})$;
  Score generated trajectories with frozen $\phi$}: $\hat{y}_{\phi}(\tau)$;
  Update $\theta$ by maximizing Eq. 3.2.2.
end while
Algorithm 1 AIGB-Pearl (Planning with EvaluAtor via RL)
Appendix E Additional Experiments
E.1 Simulated Experiment Settings
We include the detailed simulated experiment settings in Table 5. Specifically, we consider the bidding process in a day, where the bidding episode is divided into 96 time steps. Thus, the duration between two adjacent time steps $t$ and $t+1$ is 15 minutes. The number of impression opportunities between time steps $t$ and $t+1$ fluctuates from 100 to 500. The maximum and minimum budgets of advertisers are 1000 Yuan and 4000 Yuan, respectively. The upper bound of the bid price is 1000 Yuan, and the value of impressions ranges from 0 to 1.
Hardware Resource. The simulated experiments are conducted based on an NVIDIA T4 Tensor Core GPU. We use 10 CPUs and 200G memory.
Table 5: Settings of the simulated experiments.
Parameters Values
Number of advertisers 30
Time steps in an episode, $T$ 96
Minimum number of impressions within a time step 50
Maximum number of impression within a time step 300
Minimum budget 1000 Yuan
Maximum budget 4000 Yuan
[…]
Minimum bid price, $a_{\text{min}}$ 0 Yuan
Maximum bid price, $a_{\text{max}}$ 1000 Yuan
Maximum market price, $p_{M}$ 1000 Yuan
Table 6: Settings of the real-world experiments.
Parameters Values
Number of advertisers 6,000
Time steps in an episode, $T$ 96
Minimum number of impression within a time step 100
Maximum number of impressions within a time step 2,500
Minimum budget 50 Yuan
Maximum budget 10,000 Yuan
[…]
Minimum bid price, $a_{\text{min}}$ 0 Yuan
Maximum bid price, $a_{\text{max}}$ 1000 Yuan
Maximum market price, $p_{M}$ 1000 Yuan
E.2 Real-world Experiment Settings
We include the detailed real-world experiment settings in Table 6. Specifically, we consider the bidding process in a day, where the bidding episode is divided into 96 time steps. Thus, the duration between two adjacent time steps $t$ and $t+1$ is 15 minutes. The number of impression opportunities between time steps $t$ and $t+1$ fluctuates from 100 to 2,500. The maximum and minimum budgets of advertisers are 50 Yuan and 10,000 Yuan, respectively. The upper bound of the bid price is 1000 Yuan, and the value of impressions ranges from 0 to 1.
Hardware Resource. The training process in the real-world experiments is conducted using 10 NVIDIA T4 Tensor Core GPUs in a distributed manner. For each distributional worker, we use 10 CPUs and 200 GB of memory.
E.3 Real-world Experiments on TargetROAS Bidding Problem
In addition to the budget-constrained auto-bidding problem, we also apply the proposed AIGB-Pearl algorithm to a more challenging type of auto-bidding problem, named TargetROAS, with an extra ROI (Return on Investment) constraint. We conduct a real-world experiment on TaoBao, with 300k advertisers over 22 days, evaluating the performance of AIGB-Pearl on the TargetROAS auto-bidding problem. The results are given in table 7. AIGB-Pearl achieves a +5.1% improvement in GMV compared to the SOTA auto-bidding method, DiffBid, demonstrating its effectiveness in managing more complex and realistic constraints.
Table 7: Overall performance of TargetROAS in real-world A/B test, involving 300k advertisers over 22 days.
Methods GMV BuyCnt ROI Cost
DiffBid 779,642,891 11,519,082 4.68 166,544,918
AIGB-Pearl (ours) 819,550,812 11,886,501 4.70 174,234,673
$\bm{\Delta}$ +5.1% +3.2% +0.5% +4.6%
E.4 Pathological Trajectory Behavior Explanation
In industrial practice, stable and effective metrics have been developed to evaluate pathological behaviors. For the case of the budget-constrained auto-bidding problem with bidding cycles structured as 24-hour episodes (T = 96 time steps), the following three key metrics are commonly used to identify pathological behaviors:
  • Excessive budget consumption: there exists a time step $t$ such that the cost between time step $t$ and $t+1$ exceeds $10\%$ of the budget $B$;
  • Forward- (or Backward-) trending pacing: the cost between time step $1$ and $24$ (or between time step T24T-24 and $T$) exceeds $40\%$ (or $40\%$) of the budget $B$;
  • Underutilization of available budgets: the total cost over the bidding episode is lower than $90\%$ of the budget $B$.
E.5 Training Stability
We present additional comparisons between the training curves of the offline RL with bootstrapping and those of AIGB-Pearl in Fig. 6, Fig. 7, Fig. 8, and Fig. 9 concerning:
  • Cumulative Rewards: the main performance index of the considered auto-bidding problem;
  • Online Rate: the ratio between the bidding period before the budget runs out and the total bidding period. A larger Online Rate indicates a better performance.
  • Bad Case Rate: the ratio between the number of “bad” trajectories and the total number of generated trajectories. A lower Bad Case Rate indicates a better performance.
  • Cost Rate: the ratio between the cost and the budget. A larger cost rate indicates a better performance.
We can see that the offline RL method tends to suffer from significant instability throughout training, showing high variance across different seeds. In contrast, AIGB-Pearl achieves much smoother and more consistent learning progress, demonstrating the improved training stability.
Figure 6: Learning curves of cumulative rewards between offline RL with bootstrapping method and AIGB-Peral under 10 seeds.
Figure 7: Learning curves of online rate between offline RL with bootstrapping method and AIGB-Peral under 10 seeds.
Figure 8: Learning curves of bad case rate between offline RL with bootstrapping method and AIGB-Peral under 10 seeds.
Figure 9: Learning curves of cost rate between offline RL with bootstrapping method and AIGB-Peral under 10 seeds.
Appendix F LLM Usage
The authors have used Large Language Models (LLMs) exclusively for grammar checking and lexical refinement during the writing process. No LLM-generated content, data analysis, or substantive contributions to the research methodology, results, or conclusions are involved in this work.

### v3

@@
 To address this, we propose AIGB-Pearl (Planning with EvaluAtor via RL), a novel method that integrates generative planning and policy optimization.
-The core of AIGB-Pearl lies in constructing a trajectory evaluator for scoring generation quality and designing a provably sound KL-Lipschitz-constrained score maximization scheme to ensure safe and efficient generalization beyond the offline dataset.
+The core of AIGB-Pearl lies in constructing a trajectory evaluator for scoring generation quality and designing a provably sound KL-Lipschitz-constrained score maximization scheme to ensure safe and efficient exploration beyond the offline dataset.
 A practical algorithm incorporating the synchronous coupling technique is further devised to ensure the model regularity required by the proposed scheme.
@@
 Specifically, the generation quality of the generative model remains unknown during training, making it infeasible to explore new trajectories beyond the offline dataset; (ii) no dedicated offline RL algorithm exists for AIGB.
-In particular, theoretical analysis that guarantees safe generalization and mitigates OOD issues for generative models in auto-bidding remains largely unexplored.
-To address these critical challenges, we propose AIGB-Pearl (Planning with EvaluAtor via RL), an RL-enhanced version of AIGB that learns a trajectory evaluator to score generation quality and drive exploration through continuous interaction with the generative model in AIGB via RL.
-The evaluator is trained through supervised learning, leveraging the offline dataset.
-Notably, to mitigate the OOD problem and ensure reliable usage of the evaluator, we first examine the theoretical upper bound on the evaluator’s bias.
-Then, guided by this analysis, we establish a KL-Lipschitz-constrained score maximization objective with a provable sub-optimality bound, enabling safe and effective exploration for the generative model beyond the offline data.
+In particular, theoretical approaches to ensure safe generalization and mitigate OOD issues for generative models in auto-bidding remain largely unexplored.
+To address these critical challenges, we propose AIGB-Pearl (Planning with EvaluAtor via RL), an RL-enhanced version of AIGB that learns a trajectory evaluator to score generation quality and drive exploration of the generative model through continuous interaction.
+The evaluator is trained through supervised learning on the offline dataset.
+Crucially, to mitigate the OOD problem and ensure reliable evaluator utilization, we examine the theoretical upper bound on the evaluator’s bias.
+Then, guided by this analysis, we establish a KL-Lipschitz-constrained score maximization objective with a provable sub-optimality bound, enabling safe and effective exploration beyond the offline data.
 Moreover, to perform constrained score maximization, we design a practical algorithm incorporating the synchronous coupling technique, which helps ensure the Lipschitz requirement of the generative model.
-Besides, we note that AIGB-Pearl operates without the need for bootstrapping, leading to greater training stability compared to offline RL methods.
-To summarize, our contributions in this paper are fourfold: (i) we propose a novel generative auto-bidding method, AIGB-Pearl, that enables continuous improvement in generation quality through exploration beyond the offline dataset; (ii) we propose a provable KL-Lipschitz constrained score maximization objective with a sub-optimality bound, ensuring a safe and effective generalization for AIGB; (iii) we devise a practical algorithm with synchronous coupling that effectively ensures the Lipschitz requirement for the generative model; (iv) extensive simulated and real-world experiments demonstrate that AIGB-Pearl achieves SOTA performance and verify the effectiveness of the developed techniques in enhancing safe and effective generalization.
+In addition, we note that AIGB-Pearl operates without the need for bootstrapping, exhibiting greater training stability compared to offline RL methods.
+To summarize, our contributions in this paper are fourfold: (i) we propose a novel generative auto-bidding method, AIGB-Pearl, that enables continuous improvement in generation quality through exploration beyond the offline dataset; (ii) we propose a provable KL-Lipschitz constrained score maximization objective with a sub-optimality bound, ensuring a safe and effective generalization beyond the offline dataset; (iii) we devise a practical algorithm with synchronous coupling that effectively ensures the Lipschitz requirement for the generative model; (iv) extensive simulated and real-world experiments demonstrate that AIGB-Pearl achieves SOTA performance and verify the effectiveness of the developed techniques in enhancing safe and effective generalization.
 2 Preliminaries
@@
 2 results in the infamous out-of-distribution (OOD) problem (Fujimoto et al., 2019), making the policy erroneously deviate from the offline dataset $\mathcal{D}$.
-As a standard solution, offline RL (Yu et al., 2020; Kumar et al., 2020; Kidambi et al., 2020; Wang et al., 2022) constrains the policy’s behavior near $\mathcal{D}$ during TD learning, enabling reliable generalization beyond the offline dataset.
+As a standard solution, offline RL (Yu et al., 2020; Kumar et al., 2020; Kidambi et al., 2020; Wang et al., 2022b) constrains the policy’s behavior near $\mathcal{D}$ during TD learning, enabling reliable generalization beyond the offline dataset.
 However, offline RL methods notoriously suffer from training instability caused by TD-learning (Peng et al., 2024), where the bootstrapped value of the Q function serves as its training label, resulting in an erroneous ground truth.
@@
 Unlike RL methods, the AI-generated auto-bidding (AIGB) (Guo et al., 2024) treats the auto-bidding problem as a sequence generation task.
-Specifically, a generative model is employed to fit the conditional trajectory distribution $p_{\theta}(\tau|y(\tau))$ within the offline dataset $\mathcal{D}$, i.e.,
+Specifically, a conditional generative model is employed to fit the conditional trajectory distribution $p_{\theta}(\tau|y(\tau))$ within the offline dataset $\mathcal{D}$, i.e.,
 $\max_{\theta}\;\mathbb{E}_{(\tau,y(\tau))\sim\mathcal{D}}[\log p_{\theta}(\tau|y(\tau))],$ (3)
@@
 Let $y_{m}>0$ be the maximum trajectory quality in $\mathcal{D}$, we have $\forall y\in\mathcal{D},y\in[0,y_{m}]$.
-During inference, AIGB follows a planning-and-control architecture.
-Specifically, at each time step, a trajectory is sampled from the trained generative model that acts as the planner, with a manually set condition $y^{*}\triangleq(1+\epsilon)y_{m}$, where $\epsilon>0$ is a hyper-parameter with usually a small value.
-Then, an extra off-the-shelf inverse dynamic model (Agrawal et al., 2016), acting as the controller, is employed to compute the action.
+During the inference stage, AIGB follows a planning-and-control architecture.
+Specifically, at each time step, a trajectory is sampled from the generative model, acting as the planner, with a manually set condition $y^{*}\triangleq(1+\epsilon)y_{m}$, where $\epsilon>0$ is a hyper-parameter with usually a small value.
+Then, an extra off-the-shelf inverse dynamic model (Agrawal et al., 2016), acting as the controller, is employed to derive the action.
 See Appendix B for detailed descriptions.
@@
 To this end, we propose AIGB-Pearl (Planning with EvaluAtor via RL) that constructs a trajectory evaluator (referred to as the evaluator for simplicity) to integrate RL methods into AIGB’s planner training.
-Specifically, the evaluator learns a score $\hat{y}_{\phi}(\tau)$ to estimate the trajectory quality $y(\tau)$ via supervised learning, where $\phi$ denotes the evaluator parameter.
-Then, with $\phi$ fixed, the planner tries to maximize the score of its generation through constant interactions with the evaluator, as shown in Fig.
+Specifically, the evaluator learns a score $\hat{y}_{\phi}(\tau)$ to estimate the trajectory quality $y(\tau)$ via supervised learning, i.e., $\min_{\phi}\mathbb{E}_{\tau\sim\mathcal{D}}[(\hat{y}_{\phi}(\tau)-y(\tau))^{2}]$, where $\phi$ denotes the evaluator parameter.
+Then, with $\phi$ fixed, the planner tries to maximize the score of its generation through iterative interactions with the evaluator, as shown in Fig.
 1.
@@
 where the condition is fixed to $y^{*}$ in both training and inference stages to ensure consistency.
-The effectiveness of AIGB-Pearl fundamentally hinges on the evaluator’s reliability.
-However, given the offline setting of the considered auto-bidding problem, the evaluator learning is confined to the offline dataset $\mathcal{D}$, and directly solving Eq.
-4 can induce the infamous OOD problem due to the evaluator’s generalization limits, potentially resulting in the planner’s true performance degradation.
-To address this challenge, we begin by examining the theoretical bounds on the evaluator’s bias.
-Then, guided by this analysis, we propose a KL-Lipschitz-constrained score maximization objective for the planner to ensure reliable evaluator utilization in Section 3.1.
-This objective is theoretically justified by a sub-optimality bound, as established in Section 3.1.1.
-Finally, a practical algorithm is presented in Section 3.2, where a synchronous coupling method is employed to realize the planner’s Lipschitz constraint.
+As can be observed, the effectiveness of AIGB-Pearl fundamentally hinges on the reliability of the evaluator.
+However, given the offline nature of the considered auto-bidding problem, the evaluator learning is restricted to the fixed dataset $\mathcal{D}$.
+As a result, directly solving Eq.
+4 can induce the infamous OOD problem due to the evaluator’s generalization limits, potentially degrading the planner’s true performance.
+Crucially, there remains a lack of theoretical approaches to the OOD problem in RL-based training of the generative model.
+To address this challenge, we examine the theoretical bounds on the evaluator’s bias.
+Then, guided by this analysis, we propose a KL-Lipschitz-constrained score maximization objective for the planner in Section 3.1 to ensure reliable utilization of the evaluator.
+Notably, this objective is theoretically justified by a sub-optimality bound established in Section 3.1.1.
+Finally, a practical algorithm is presented in Section 3.2, where tailored regularization methods are introduced to enforce Lipschitz continuity in both the evaluator and the planner.
 3.1 KL-Lipschitz-constrained score maximization
@@
 Our basic idea is to optimize $\theta$ within a domain where the gap between the planner’s score $L(\theta)$ and its true performance $J(\theta)\triangleq\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)]$ is bounded by a small certifiable upper bound.
-This ensures the score maximization occurs only in regions where the evaluator is reliable.
+This ensures the score maximization occurs only in regions where the evaluator is relatively reliable.
 $|J(\theta)-L(\theta)|=|\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[y(\tau)]-\mathbb{E}_{\tau\sim p_{\theta}(\tau|y^{*})}[\hat{y}_{\phi}(\tau)]|.$ (5)
@@
 The trajectory quality $y(\tau)$ is $\sqrt{T}R_{m}$-Lipschitz continuous with respect to the Frobenius norm.
-Based on this property, we incorporate a $\sqrt{T}R_{m}$-Lipschitz continuity constraint into the evaluator’s design (as described in Section 3.2.1), and denote the Lipschitz constant of the resulting evaluator as $k\sqrt{T}R_{m}$, where $k\geq 0$ reflects the violation degree.
+Based on this property, we incorporate a $\sqrt{T}R_{m}$-Lipschitz continuity constraint into the evaluator’s learning (as described in Section 3.2.1), and denote the Lipschitz constant of the resulting evaluator as $k\sqrt{T}R_{m}$, where $k\geq 0$ reflects the violation degree.
 A tighter satisfaction of the Lipschitz constraint by the evaluator results in $k$ closer to $1$.
@@
 where $W_{1}$ denotes the 1-Wasserstein distance.
-Note that $\delta_{D}$ could be regulated to a small value via supervised training of the evaluator.
-Consequently, bounding the planner’s performance gap requires constraining the following two factors:
+Note that $\delta_{D}$ could be regulated to a small value via supervised training of the evaluator, and $k$ depends on the Lipschitz property of the resulting evaluator  [^2] Note that $k$ cannot approach zero without compromising $\delta_{D}$, as excessively small $k$ prevents the evaluator from fitting the offline dataset $\mathcal{D}$..
+Consequently, bounding the evaluator bias in the planner’s performance requires constraining the following two factors:
   • the planner’s Lipschitz sensitivity to condition $y$ (the first Wasserstein term)
@@
 The proof is given in Appendix C.3.
-Therefore, we constrain the planner’s Lipschitz constant to a positive value $L_{p}>0$ to ensure the boundedness of the first Wasserstein term, where $L_{p}$ is a hyperparameter whose recommended range is provided later in Eq.
+Therefore, we constrain the planner’s Lipschitz constant to a positive value $L_{p}>0$ to ensure the boundedness of the first Wasserstein term, where $L_{p}$ is a hyperparameter whose lower bound is provided later in Eq.
 10.
@@
 Remark 1.
-Intuitively, the KL and Lipschitz constraints collectively ensure the planner’s generation under condition $y^{*}$ remains within a certified neighborhood of the offline dataset $\mathcal{D}$, while the evaluator maintains high accuracy within this $\mathcal{D}$-proximal region as it is trained with the offline dataset.
+Intuitively, the KL and Lipschitz constraints collectively ensure the planner’s generation under condition $y^{*}$ remains within a certified neighborhood of the offline dataset $\mathcal{D}$.
+At the same time, the evaluator maintains high accuracy within this $\mathcal{D}$-proximal region as it is trained on $\mathcal{D}$, and the Lipschitz regularization on the evaluator bounds its sensitivity to input perturbations, preventing drastic value fluctuations in OOD regions and promoting more reliable extrapolation.
 Therefore, Eq.
-8 can realize reliable evaluator-guided score maximization.
+8 with a Lipschitz-regularized evaluator can realize reliable evaluator-guided score maximization.
 3.1.1 sub-optimality gap bound
@@
 Theorem 3 (Sub-optimality Gap Bound).
-Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ be the expected distance between the optimal trajectory distribution and the trajectory distribution of the offline dataset $\mathcal{D}$.
+Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y^{*})\|p_{D}(\tau|y))]$ be the expected distance between the optimal trajectory distribution and the trajectory distribution of the offline dataset $\mathcal{D}$.
 The true performance gap between the optimal parameter $\theta^{*}$ and the solution $\hat{\theta}$ to Eq.
@@
 Theoretical Result Analysis.
-In Theorem 3, the constants $\delta_{M},R_{m},T,y_{m},\epsilon,\rho$ characterize domain-specific properties of the auto-bidding task and the offline dataset $\mathcal{D}$.
+In Theorem 3, the constants $\delta_{M},R_{m},T,\epsilon,y_{m},\rho$ characterize domain-specific properties of the auto-bidding task and the offline dataset $\mathcal{D}$.
 Nonetheless, a lower training error $\delta_{D}$ of the evaluator and a closer $k$ to $1$ correspond to a smaller sub-optimality gap.
@@
 where $\beta_{1}>0$ is a hyper-parameter, $\cdot]_{+}\triangleq\max\{0,\cdot\}$.
+Moreover, to further enhance the accuracy of the evaluator, we devise two specific techniques, including the LLM Embedding enhancement and pair-wise learning, whose details are given in Appendix D.1.
 3.2.2 Lipschitz Planner With Synchronous Coupling
@@
 Wasserstein Upper Bound as Surrogate.
-Accurate computation of this Wasserstein distance term is challenging, as it requires finding the optimal coupling between $p_{\theta}(\tau|y_{1})$ and $p_{\theta}(\tau|y_{2})$ that minimizes the expected transportation cost.
+Accurate calculation of this Wasserstein distance term is computationally challenging, as it requires finding the optimal coupling between $p_{\theta}(\tau|y_{1})$ and $p_{\theta}(\tau|y_{2})$ that minimizes the expected transportation cost.
 Nonetheless, we can choose a certain coupling $\gamma\in\Gamma(p_{\theta}(\tau|y_{1}),p_{\theta}(\tau|y_{2}))$ to obtain an upper bound of this Wasserstein term, i.e.,
@@
 Moreover, if we fix the predicted variance $\sigma_{\theta}$ of the planner as a fixed constant, then the expression of $\hat{W}_{1}(y_{1},y_{2};\theta)$ can be further simplified to $\hat{W}_{1}(y_{1},y_{2};\theta)=\sum_{t}\|\mu_{\theta}(s^{1}_{1:t},y_{1},t)-\mu_{\theta}(s^{2}_{1:t},y_{2},t)\|$.
-The overall AIGB-Pearl algorithm is summarized in Algorithm 1 in Appendix D due to page limits.
+The overall AIGB-Pearl is summarized in Algorithm 1 in Appendix D.2 due to page limits.
+Remark 2.
+We note that AIGB-Pearl can fully leverage the generalization capability of the conditional generative model.
+Specifically, the model acquires an implicit understanding of how trajectory characteristics vary with $y$ through conditional behavior cloning on $\mathcal{D}$.
+This enables the model not only to learn a rich and well-calibrated distribution over trajectories for conditions $y\in\mathcal{D}$, but also to generalize smoothly beyond the observed range, providing a high-quality initialization for OOD condition $y^{*}$.
+(In practice, we often initialize the search at $y^{*}$ using a pre-trained conditional generator as described in Algorithm 1.) Subsequently, under the guidance of reward maximization and constrained by Lipschitz continuity with respect to $y$, the model safely and efficiently explores improved trajectories without collapsing into unrealistic or unstable solutions.
+Figure 1: AIGB-Pearl (Planning with EvaluAtor via RL) constructs a trajectory evaluator to score the trajectory quality and let the planner maximize the obtained score under the KL-Lipschitz constraint through continuous interaction with the evaluator.
+A synchronous coupling method is used to estimate the Wasserstein term in the Lipschitz penalty.
+Table 1: Overall performance (GMV) in simulated experiments with 30 advertisers.
+$\Delta$ indicates the relative improvement of AIGB-Pearl against the most competitive baseline (which is underlined).
+Note that the absolute values are normalized without specific meanings; only $\Delta$ matters.
+Budget USCB BCQ CQL IQL Diff-QL MOPO DT DiffBid AIGB-Pearl $\bm{\Delta}$
+1.5k 454.25 454.72 461.82 456.80 469.73 470.38 477.39 480.76 502.98 +4.62%
+2.0k 482.67 483.50 475.78 486.56 487.91 489.27 507.30 511.17 521.84 +2.09%
+2.5k 497.66 498.77 481.37 518.27 510.83 523.91 527.88 531.29 545.03 +2.59%
+3.0k 500.60 501.86 491.36 549.19 552.73 549.01 550.66 556.32 574.17 +3.21%
+Table 2: Overall performance in real-world A/B tests, involving 6k advertisers over 19 days.
+Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DiffBid 76,390,174 650,962 5.31 14,395,290 USCB 52,182,805 516,994 4.92 10,598,486 AIGB-Pearl 78,676,009 665,173 5.41 14,551,054 AIGB-Pearl 53,973,101 520,796 5.13 10,515,772 $\bm{\Delta}$ +3.00% +2.20% +1.89% +1.10% $\bm{\Delta}$ +3.43% +0.74% +4.24% -0.78% Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DT 34,808,665 341,995 5.61 6,205,665 MOPO 51,674,071 579,332 3.08 16,771,892 AIGB-Pearl 35,957,933 344,194 5.77 6,246,512 AIGB-Pearl 53,292,945 591,741 3.23 16,475,670 $\bm{\Delta}$ +3.30% +0.64% +0.16% +0.66% $\bm{\Delta}$ +3.13% +2.14% +4.87% -1.77%
 4 Experiments
@@
 (Section 4.4).
-The training stability of AIGB-Pearl is studied in Appendix E.5.
-Figure 1: AIGB-Pearl (Planning with EvaluAtor via RL) constructs a trajectory evaluator to score the trajectory quality and let the planner maximize the obtained score under the KL-Lipschitz constraint through continuous interaction with the evaluator.
-A synchronous coupling method is used to estimate the Wasserstein term in the Lipschitz penalty.
+Moreover, the training stability of the planner is analyzed in Appendix E.5, and the accuracy of the evaluator is given in Appendix E.6, respectively.
 4.1 Experiment Setup
@@
 We compare our method with the state-of-the-art AIGB methods, including DiffBid (Guo et al., 2024) and DT (Chen et al., 2021a) that learn from conditional behavior cloning of the offline dataset with a diffusion model and a Causal Transformer, respectively.
-We also compare our method with RL auto-bidding methods, including USCB (He et al., 2021) that learns the auto-bidding policy in a manually constructed advertising system with DDPG (Silver et al., 2014); and offline RL auto-bidding methods, including model-free offline RL methods BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022) and Diff-QL (Wang et al., 2022), and model-based offline RL method MOPO (Yu et al., 2020).
+We also compare our method with RL auto-bidding methods, including USCB (He et al., 2021) that learns the auto-bidding policy in a manually constructed advertising system with DDPG (Silver et al., 2014); and offline RL auto-bidding methods, including model-free offline RL methods BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022) and Diff-QL (Wang et al., 2022b), and model-based offline RL method MOPO (Yu et al., 2020).
 Performance Index.
@@
 Note that larger values of GMV, BuyCnt, and ROI with a Cost oscillating within an acceptable tolerance ($\pm 2\%$) indicate a better performance.
-4.2 Overall Performance
-Table 1: Overall performance (GMV) in simulated experiments with 30 advertisers.
-$\Delta$ indicates the relative improvement of AIGB-Pearl against the most competitive baseline (which is underlined).
-Note that the absolute values are normalized without specific meanings; only $\Delta$ matters.
-Budget USCB BCQ CQL IQL Diff-QL MOPO DT DiffBid AIGB-Pearl $\bm{\Delta}$
-1.5k 454.25 454.72 461.82 456.80 469.73 470.38 477.39 480.76 502.98 +4.62%
-2.0k 482.67 483.50 475.78 486.56 487.91 489.27 507.30 511.17 521.84 +2.09%
-2.5k 497.66 498.77 481.37 518.27 510.83 523.91 527.88 531.29 545.03 +2.59%
-3.0k 500.60 501.86 491.36 549.19 552.73 549.01 550.66 556.32 574.17 +3.21%
-Table 2: Overall performance in real-world A/B tests, involving 6k advertisers over 19 days.
-Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DiffBid 76,390,174 650,962 5.31 14,395,290 USCB 52,182,805 516,994 4.92 10,598,486 AIGB-Pearl 78,676,009 665,173 5.41 14,551,054 AIGB-Pearl 53,973,101 520,796 5.13 10,515,772 $\bm{\Delta}$ +3.00% +2.20% +1.89% +1.10% $\bm{\Delta}$ +3.43% +0.74% +4.24% -0.78% Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DT 34,808,665 341,995 5.61 6,205,665 MOPO 51,674,071 579,332 3.08 16,771,892 AIGB-Pearl 35,957,933 344,194 5.77 6,246,512 AIGB-Pearl 53,292,945 591,741 3.23 16,475,670 $\bm{\Delta}$ +3.30% +0.64% +0.16% +0.66% $\bm{\Delta}$ +3.13% +2.14% +4.87% -1.77%
 Table 3: Generalization performance in real-world A/B tests with unseen advertisers against AIGB methods, involving 4k advertisers over 19 days.
 Methods GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost DiffBid 67,092,973 553,020 5.39 12,444,306 DT 30,562,007 300,271 5.61 5,450,573 AIGB-Pearl 69,252,539 565,776 5.53 12,534,379 AIGB-Pearl 31,502,309 305,202 5.74 5,484,473 $\bm{\Delta}$ +3.32% +2.31% +2.48% +0.72% $\bm{\Delta}$ +3.08% +1.64% +2.32% +0.62%
+Figure 2: Trajectory Generation Visualization.
+Three cases are presented.
+Here, the AIGB-Pearl generates plausible trajectories, whereas its variant without the KL-Lipschitz constraint produces generations that significantly deviate from the reference and exhibit evident issues.
+Table 4: Ablation Study. The effectiveness of the KL constraint and the Lipschitz constraint in Real-world A/B tests, involving 6k advertisers over 8 days.
+AIGB-Pearl GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost w/o KL 30,906,963 292,605 4.25 7,269,018 w/o Lipschitz 32,284,972 268,551 5.73 5,634,304 with KL 31,243,688 292,783 4.26 7,342,485 with Lipschitz 32,869,329 281,979 5.79 5,678,252 $\bm{\Delta}$ +1.09% +0.06% +0.08% +1.01% $\bm{\Delta}$ +1.81% +0.50% +1.05% +0.78%
+4.2 Overall Performance
 To answer RQ(1): Table 1 shows that our method consistently outperforms all baselines in GMV across all four budget levels in simulated experiments.
@@
 It can be seen that the KL constraint contributes +1.1% improvement in GMV, and the Lipschitz constraint provides +1.8% improvement in GMV, demonstrating their respective roles in enhancing AIGB-Pearl’s performance.
-Table 4: Ablation Study. The effectiveness of the KL constraint and the Lipschitz constraint in Real-world A/B tests, involving 6k advertisers over 8 days.
-AIGB-Pearl GMV BuyCnt ROI Cost Methods GMV BuyCnt ROI Cost w/o KL 30,906,963 292,605 4.25 7,269,018 w/o Lipschitz 32,284,972 268,551 5.73 5,634,304 with KL 31,243,688 292,783 4.26 7,342,485 with Lipschitz 32,869,329 281,979 5.79 5,678,252 $\bm{\Delta}$ +1.09% +0.06% +0.08% +1.01% $\bm{\Delta}$ +1.81% +0.50% +1.05% +0.78%
 Visualization.
@@
 and exhibit clear pathological behaviors—such as excessive budget consumption, backward-trending pacing, and underutilization of available budgets (see Appendix E.4 for explanation)—which further validate the KL-Lipschitz constraint necessity.
-Figure 2: Trajectory Generation Visualization.
-Three cases are presented.
-Here, the AIGB-Pearl generates plausible trajectories, whereas its variant without the KL-Lipschitz constraint produces generations that significantly deviate from the reference and exhibit evident issues.
 4.4 Lipschitz Value Examination
@@
 These methods allow policy learning from pre-collected datasets without requiring online interaction.
-Moreover, offline RL, such as Diffusion-QL (Wang et al., 2022), adopts generative models as the policy model architecture for better expressive capacity.
+Moreover, offline RL, such as Diffusion-QL (Wang et al., 2022b), adopts generative models as the policy model architecture for better expressive capacity.
 However, RL-based methods often suffer from training instability caused by bootstrapping and alternating training paradigms between critics and actors.
@@
 $|J(\theta^{*})-L(\theta^{*})|$ $\leq\mathbb{E}_{y\sim p_{D}(y)}\mathbb{E}_{\tau\sim p_{D}(\tau|y)}f(\tau)+\mathbb{E}_{y\sim p_{D}(y)}\bigg[\mathbb{E}_{\tau\sim p_{\theta^{*}}(\tau|y^{*})}f(\tau)-\mathbb{E}_{\tau\sim p_{D}(\tau|y)}f(\tau)\bigg]$
-$\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))],$ (41)
+$\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\mathbb{E}_{y\sim p_{D}(y)}\bigg[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))\bigg],$ (41)
 where $W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))$ denotes the probability distribution distance between the optimal planner and the offline dataset.
@@
 C.4, we have:
-$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))]\leq\sqrt{\frac{2}{\rho}\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}(\tau|y)}\|p_{D}(\tau|y))]}$ (42)
-Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y)\|p_{D}(\tau|y))]$ be the distance between the optimal trajectory distribution and the offline dataset trajectory distribution.
+$\mathbb{E}_{y\sim p_{D}(y)}[W_{1}(p_{\theta^{*}}(\tau|y^{*}),p_{D}(\tau|y))]\leq\sqrt{\frac{2}{\rho}\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y^{*})\|p_{D}(\tau|y))]}$ (42)
+Let $\delta_{M}\triangleq\mathbb{E}_{y\sim p_{D}(y)}[D_{KL}(p_{\theta^{*}}(\tau|y^{*})\|p_{D}(\tau|y))]$ be the distance between the optimal trajectory distribution and the offline dataset trajectory distribution.
 We have:
@@
 (3) Evaluator Bias in $p_{\hat{\theta}}$. Since $\hat{\theta}$ satisfies the KL and Lipschitz constraint in Eq. 8, we can use the results in Theorem 2, Eq. 6 and Eq. 7 to obtain:
-$|J(\hat{\theta})-L(\hat{\theta})|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}],$ (45)
-Overall, combining the results in Eq. 43, Eq. C.5, and Eq. 45, we have:
-$J(\theta^{*})-J(\hat{\theta})\leq 2\delta_{D}+(1+2k)\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+\sqrt{2\delta_{K}/\rho}+(1+\epsilon)y_{m}L_{p}\bigg].$ (46)
+$|J(\hat{\theta})-L(\hat{\theta})|\leq\delta_{D}+(1+k)\sqrt{T}R_{m}\bigg[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg],$ (45)
+Overall, combining the results in Eq. C.5, Eq. 43, Eq. C.5, and Eq. 45, we have:
+[…]
+$\quad+k\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg]$
+$\quad+\delta_{D}+(1+k)\sqrt{T}R_{m}\bigg[(1+\epsilon)y_{m}L_{p}+\sqrt{2\delta_{K}/\rho}\bigg]$
+$=2\delta_{D}+(1+2k)\sqrt{T}R_{m}\bigg[\sqrt{2\delta_{M}/\rho}+\sqrt{2\delta_{K}/\rho}+(1+\epsilon)y_{m}L_{p}\bigg].$ (46)
 This concludes the proof.
@@
 $p_{\theta}(s_{1:T}|y)=\Pi_{t}p_{\theta}(s_{t}|s_{1:t-1},y).$ (47)
-Then, we have:
+Then, we can derive the gradient as follows:
 $\nabla_{\theta}L(\theta)$ $=\nabla_{\theta}\int_{\tau}p_{\theta}(\tau|y^{*})\hat{y}_{\phi}(\tau)\mathrm{d}\tau$
@@
 $=\mathbb{E}_{s_{1:T}\sim p_{\theta}(s_{1:T}|y^{*})}\bigg[\sum_{t}\nabla_{\theta}\log p_{\theta}(s_{t}|s_{1:t-1},y^{*})\hat{y}_{\phi}(\tau)\bigg].$ (48)
-Appendix D AIGB-Pearl Algorithm Summary
-Algorithm 1 summarizes the training process of AIGB-Pearl. The development of AIGB-Pearl is supported by ROLL (Wang et al., 2025).
-Input : Offline dataset $\mathcal{D}$, desired condition $y^{*}$, hyper-parameters $\beta_{1}$, $\beta_{2},\beta_{3}$.
+In conclusion, we have:
+$\nabla_{\theta}L(\theta)=\mathbb{E}_{s_{1:T}\sim p_{\theta}(s_{1:T}|y^{*})}\bigg[\sum_{t}\nabla_{\theta}\log p_{\theta}(s_{t}|s_{1:t-1},y^{*})\hat{y}_{\phi}(\tau)\bigg].$ (49)
+Appendix D AIGB-Pearl Practical Algorithm
+D.1 Additional Designs For Evaluator Accuracy Enhancement
+To further enhance the reliability of the trajectory evaluator, we design two additional techniques for the evaluator learning.
+Specifically, as described in the following, we (i) integrate an LLM embedding into its input feature for better representational capacity; and (ii) employ the pair-wise loss for better score estimation accuracy.
+Fig.
+6 illustrates the complete evaluator learning losses.
+The effectiveness of these two methods is studied in Appendix E.6.
+D.1.1 LLM Embedding Enhancement
+Motivated by the success of integrating user-specific features into the recommendation system (Chen et al., 2015), we incorporate advertiser-specific features into the trajectory evaluator to enhance its representational capacity for better scoring accuracy.
+Note that some advertiser-specific features are textual in nature—such as product titles, categories, and reviews—and are therefore difficult to incorporate directly into the vectorized trajectory $\tau$.
+To address this, we construct a prompt containing all such textual attributes and employ a pre-trained large language model (LLM), QWen2.5-1.5B-Instruct, with general world knowledge to extract $T$ representation, which we refer to as the LLM embedding.
+Specifically, we use the output of the last hidden layer as the LLM embedding.
+This embedding is then used as an additional positional encoding in the evaluator.
+The prompt template is given below:
+LLM Prompt Template.
+I am an [advertising platform] advertiser, operating the [brand name] brand in the category of [category name], classified as a [advertiser tier] tier advertiser.
+I have a product titled [product name] currently running in advertising campaigns.
+This product belongs to the leaf category of [leaf category], with a price of [product price], positioned in the [price range] price range.
+Its price ranks within the top [price ranking in the leaf category] % in the leaf category.
+Historical Average Performance: The product generates an average of [average daily transactions] daily transactions, with a GMV of [average daily GMV] driven by advertising.
+It receives an average [average daily impressions] daily impressions from search and recommendation traffic, [click numbers] clicks, [average daily BuyCnt] BuyCnt, and a GMV of [GMV].
+It ranks within the top [sales volume ranking in leaf category] % in sales volume in the leaf category, with an average transaction value of [average transaction value].
+Historical Time-based Average Performance: This product has undergone continuous advertisement exposure for number of advertising days days.
+The average hourly advertising spend distribution per day (from 0:$0$ to 24:$0$) is historical spend distribution sequence.
+The average GMV distribution across its category during this period is historical GMV distribution sequence.
+The average daily advertising budget is daily advertising budget.
+D.1.2 Pair-wise Loss
+Unlike human feedback scores used in LLM post-training (Christiano et al., 2017)—which can suffer from subjective biases in their absolute valuations—the trajectory quality $y(\tau)$ has real physical meaning and is comparable among different trajectories.
+Therefore, we can adopt a hybrid point-wise and pair-wise loss for the score to capture the absolute value of $y(\tau)$ and their relative preference, respectively.
+This approach has demonstrated superior performance in the recommendation domain (Cao et al., 2007; Lei et al., 2017; Wang et al., 2022a).
+Specifically, the pair-wise loss function can be expressed as:
+$\mathbb{E}_{(\tau_{w},\tau_{l})\sim\mathcal{D}_{p}}\bigg[\log\sigma(\hat{y}_{\phi}(\tau_{w})-\hat{y}_{\phi}(\tau_{l}))\bigg],$ (50)
+where the pair-wise loss is implemented by the typical Bradley-Terry (BT) function (Bradley & Terry, 1952).
+Here, $\mathcal{D}_{p}=\{(\tau_{w},\tau_{l},y(\tau_{w}),y(\tau_{l}))\}$ denotes a pair-wise dataset extracted from the offline dataset $\mathcal{D}$, where $\tau_{w}$ denotes the trajectory with higher trajectory quality, i.e., $y(\tau_{w})\geq y(\tau_{l})$.
+D.1.3 Overall Evaluator Loss
+Combining the previous techniques, the overall evaluator loss is:
+$l_{e}(\phi)$ $=\underbrace{\mathbb{E}_{\tau\sim\mathcal{D}}\bigg[(\hat{y}_{\phi}(\tau)-y(\tau))^{2}\bigg]}_{\text{point-wise loss}}+\beta_{4}\underbrace{\mathbb{E}_{(\tau_{w},\tau_{l})\sim\mathcal{D}_{p}}\bigg[\log\sigma(\hat{y}_{\phi}(\tau_{w})-\hat{y}_{\phi}(\tau_{l}))\bigg]}_{\text{pair-wise loss}}$
+$+\beta_{1}\underbrace{\mathbb{E}_{\tau_{1},\tau_{2}}\bigg[|\hat{y}_{\phi}(\tau_{1})-\hat{y}_{\phi}(\tau_{2})|-\sqrt{T}R_{m}\left\|\tau_{1}-\tau_{2}\right\|_{F}\bigg]_{+}}_{\text{Lipschitz penalty}},$ (51)
+where $\beta_{4}>0$ is a hyper-parameter.
+Figure 6: AIGB-Pearl Evaluator Learning.
+The evaluator loss is composed of three parts, including the point-wise loss, the pair-wise loss, and the Lipschitz penalty.
+The input representation of the evaluator is augmented with the LLM embedding to incorporate semantic information from textual advertiser features.
+D.2 Algorithm Summary
+Algorithm 1 summarizes the training process of AIGB-Pearl.
+Specifically, we calculate the Lipschitz constants of the $y(\tau)$ and $p_{D}(\tau|y)$ using the offline dataset $\mathcal{D}$, and accordingly, determine the Lipschitz constraint values for the evaluator and the planner, respectively.
+Then, we consecutively perform evaluator learning, planner pre-training, and KL-Lipschitz constrained score maximization for the planner.
+The development of AIGB-Pearl is supported by ROLL (Wang et al., 2025).
+Input : Offline dataset $\mathcal{D}$, desired condition $y^{*}$, hyper-parameters $\beta_{1}$, $\beta_{2},\beta_{3},\beta_{4}$.
 Output : Optimized $\theta$} and $\phi$
 Initialization : Randomly initialized planner parameter $\theta$}, trajectory evaluator parameters $\phi$}
-// Determining the Lipschitz Value
+// Determine the Lipschitz Constant
 Calculate the Lipschitz value of $y(\tau)$ and $p_{D}(\tau|y)$ using the offline dataset $\mathcal{D}$.
-Set the Lipschitz constraint value $L_{e}$ for the evaluator and $L_{p}$ for the planner to be bigger than the Lipschitz value of $y(\tau)$ and $p_{D}(\tau|y)$, respectively.
-// Training the trajectory evaluator
+Set the Lipschitz constraint value $\sqrt{T}R_{m}$ for the evaluator and $L_{p}$ for the planner, respectively.
+// Learn the Trajectory Evaluator
 while not converged do
@@
 end while
-// Training the generative planner
+// Pre-Train the Generative Planner
+while not converged do
+  Update $\theta$ by minimizing Eq. 3;
+end while
+// KL-Lipschitz Constrained Score Maximization
 Warm start with pretrained planner $p_{\theta}$;
@@
   Score generated trajectories with frozen $\phi$}: $\hat{y}_{\phi}(\tau)$;
-  Update $\theta$ by maximizing Eq. 3.2.2.
+  Update $\theta$ by minimizing Eq. 3.2.2.
 end while
@@
 AIGB-Pearl achieves a +5.1% improvement in GMV compared to the SOTA auto-bidding method, DiffBid, demonstrating its effectiveness in managing more complex and realistic constraints.
-Table 7: Overall performance of TargetROAS in real-world A/B test, involving 300k advertisers over 22 days.
+Table 7: The overall performance of AIGB-Pearl in TargetROAS problem in real-world A/B test, involving 300k advertisers over 22 days.
 Methods GMV BuyCnt ROI Cost
@@
   • Underutilization of available budgets: the total cost over the bidding episode is lower than $90\%$ of the budget $B$.
-E.5 Training Stability
-We present additional comparisons between the training curves of the offline RL with bootstrapping and those of AIGB-Pearl in Fig. 6, Fig. 7, Fig. 8, and Fig. 9 concerning:
+E.5 Planner Training Stability
+We present additional comparisons between the planner’s training curves of the offline RL with bootstrapping and those of AIGB-Pearl in Fig. 8, Fig. 9, Fig. 10, and Fig. 11 concerning:
   • Cumulative Rewards: the main performance index of the considered auto-bidding problem;
@@
   • Cost Rate: the ratio between the cost and the budget. A larger cost rate indicates a better performance.
-We can see that the offline RL method tends to suffer from significant instability throughout training, showing high variance across different seeds.
+We can observe that the offline RL method tends to exhibit significant instability throughout training, displaying high variance across different seeds.
 In contrast, AIGB-Pearl achieves much smoother and more consistent learning progress, demonstrating the improved training stability.
+E.6 Evaluator Accuracy
+Accuracy Metric.
+We evaluate the accuracy of the evaluator along two dimensions, including the absolute accuracy, reflecting how close the predicted scores are to ground truth qualities, and the order accuracy, reflecting the correctness of relative rankings between trajectory pairs.
+Specifically, we use the symmetric mean absolute percentage error, SMAPE, as the metric for the absolute accuracy and the AUC, defined as the ratio of correctly predicted ordinal pairs to the total number of pairs, as the metric for the order accuracy.
+The SMAPE ranges from $0\%$ to $200\%$ and AUC from $0\%$ to $100\%$.
+Lower SMAPE and larger AUC indicate better evaluator accuracy.
+In the following, we investigate the effectiveness of employing LLM embedding and pair-wise loss in the evaluator learning.
+LLM Embedding Effectiveness.
+We examine the accuracy of the trajectory evaluator without using LLM embeddings, and the results are reported in the lines of “w/o LLM” in Table 9 and Table 9.
+We can see that LLM embeddings can help improve both absolute and order accuracy.
+Fig.
+7 compares the training progress with and without LLM embeddings in terms of the SMAPE metric.
+As illustrated, the evaluator incorporating LLM embeddings exhibits faster convergence and achieves lower SMAPE compared to the one without LLM embeddings.
+The performance gain stems from the ability of LLM embeddings to encode high-level semantic information, facilitating a more nuanced understanding of the dependencies between sequential bidding states.
+Figure 7: The SMAPE training curves of the trajectory evaluator with and without LLM Embeddings. Incorporating LLM embeddings helps to achieve faster convergence and improved absolute accuracy.
+Hybrid Point-wise and Pair-wise Losses Effectiveness.
+Table 9 and Table 9 present the SMAPE and AUC of the trajectory evaluator when trained with point-wise loss only, pair-wise loss only, and a combination of both.
+It can also be seen that using only pair-wise loss results in significantly worse SMAPE performance despite some improvement in AUC.
+This suggests that while pair-wise loss can enhance ranking consistency, it falls short in providing accurate absolute value predictions.
+When both point-wise and pair-wise losses are used together, the evaluator achieves lower SMAPE and higher AUC.
+This indicates that combining these two types of losses not only improves the absolute accuracy but also enhances the order accuracy in trajectory evaluation.
+Table 8: SMAPE results from ablation experiments on the trajectory evaluator.
+SMAPE $\downarrow$ Point-wise Pair-wise Both
+w/o LLM 43.55% 196.38% 35.00%
+with LLM 38.20% 196.00% 31.60%
+Table 9: AUC results from ablation experiments on the trajectory evaluator.
+AUC $\uparrow$ Point-wise Pair-wise Both
+w/o LLM 61.57% 64.91% 70.91%
+with LLM 65.00% 71.30% 75.10%
 Figure 8: Learning curves of cumulative rewards between offline RL with bootstrapping method and AIGB-Peral under 10 seeds.